### v2

# Agnostic Smoothed Online Learning

Abstract
Classical results in statistical learning typically consider two extreme data-generating models: i.i.d. instances from an unknown distribution, or fully adversarial instances, often much more challenging statistically. To bridge the gap between these models, recent work introduced the smoothed framework, in which at each iteration an adversary generates instances from a distribution constrained to have density bounded by $\sigma^{-1}$ compared to some fixed base measure $\mu$. This framework interpolates between the i.i.d. and adversarial cases, depending on the value of $\sigma$. For the classical online prediction problem, most prior results in smoothed online learning rely on the arguably strong assumption that the base measure $\mu$ is known to the learner, contrasting with standard settings in the PAC learning or consistency literature. We consider the general agnostic problem in which the base measure is unknown and values are arbitrary. Along this direction, [15] showed that empirical risk minimization has sublinear regret under the well-specified assumption. We propose an algorithm R-Cover based on recursive coverings which is the first to guarantee sublinear regret for agnostic smoothed online learning without prior knowledge of $\mu$. For classification, we prove that R-Cover has adaptive regret $\tilde{\mathcal{O}}(\sqrt{dT/\sigma})$ for function classes with VC dimension $d$, which is optimal up to logarithmic factors. For regression, we establish that R-Cover has sublinear oblivious regret for function classes with polynomial fat-shattering dimension growth.
1 Introduction
We study the classical prediction problem in which a learner sequentially observes an instance $x_{t}\in\mathcal{X}$ and makes a prediction about a value $y_{t}\in\mathcal{Y}$ before observing the true value. The learner’s objective is to minimize the error of its predictions $\hat{y}_{t}$ compared to the true value $y_{t}$, given by some known loss function. We focus on both classification with $\mathcal{Y}=\{0,1\}$ and regression with $\mathcal{Y}=[0,1]$, but for ease of presentation the present discussion mostly concerns classification. A major question in statistical learning theory is to understand under which assumptions on the data generating process and in particular on the process generating instances $(x_{t})_{t\geq 1}$, can one give learning guarantees in the sense that the learner incurs low excess loss compared to some benchmark function class $\mathcal{F}$. Most of the literature focused on either of the two following settings.
On one extreme, one can consider that the sequence $(x_{t})_{t\geq 1}$ is fully adversarial and may depend on the actions of the learner. In this case, classical results [38, 5] show that the best one can hope for is to achieve low excess loss compared to function classes with finite Littlestone dimension. This is quite restrictive, for instance, this precludes positive results even for the simple function class of threshold functions $x\mapsto\mathbbm{1}_{x\geq x_{0}}$ on $\mathcal{X}=[0,1]$.
On the other hand, one can suppose that the instance sequence $(x_{t})_{t\geq 1}$ is i.i.d. typically under some unknown distribution $\mu$. In the PAC learning setting [49, 50, 48], one can instead ensure low excess error compared to function classes with finite VC dimension (see Definition 2) which is significantly weaker than having finite Littlestone dimension. For instance, this covers the class of linear separators for say $\mathcal{X}=\mathbb{R}^{d}$ for $d\geq 1$. In regression, this can be replaced with the notion of fat-shattering dimension (see Definition 3) [13, 37], which is a scale-dependent version of the VC dimension. In fact, when the data generating process is i.i.d. one can achieve consistency—vanishing average excess loss—without further function class assumptions [^1] Note that this differs from the PAC learning setting in the sense that guarantees are asymptotic.. For instance, in classification and when the instance space $\mathcal{X}$ is Euclidean, the simple $k$-nearest neighbor algorithm is already consistent [21, 22, 27] under reasonable choices of $k(t)$. Similar consistency results can also be achieved for general spaces [34, 29].
Ideally, one would aim to obtain similar guarantees as for the more amenable i.i.d. case under weaker statistical assumptions. Notably, there has been significant work to establish consistency results under non-i.i.d. instance processes $(x_{t})_{t\geq 1}$, including relaxations of the i.i.d. assumption such as stationary ergodic processes [42, 28, 27] or processes satisfying some form of law of large numbers [26, 47]. More recently, [31] initiated a line of work on universal learning to characterize minimal assumptions on instance processes $(x_{t})_{t\geq 1}$ for consistency [3, 10, 9, 7, 6]. These consistency results are however mostly asymptotic in nature.
Smoothed online learning.
To interpolate between the adversarial and i.i.d. case while preserving quantitative convergence rates, [43] introduced the setting of smoothed online learning. In this setting, one supposes that the process $(x_{t})_{t\geq 1}$ is generated from some limited adversary that samples $x_{t}\sim\mu_{t}$ according to some distribution $\mu_{t}$ conditional on the history, constrained to have density bounded by $1/\sigma$ with respect to some fixed distribution $\mu$ (see Definition 1). Here, $\sigma\in[0,1]$ is a parameter quantifying the smoothness of the adversary. Effectively this corresponds to a setting where the instances chosen by the adversary do not put too much mass on regions with low $\mu$-probability, which restricts the power of the adversary to explore unrelated regions of the space. Depending on the smoothness parameter $\sigma$, smoothed online learning interpolates exactly between the adversarial setting ($\sigma=0$) and the i.i.d. setting ($\sigma=1$). Recent works showed that many of the positive results from the i.i.d. case can be achieved under smooth adversaries up to paying a reasonable price in the smoothness constraint $1/\sigma$, covering a wide variety of settings from standard classification and regression [43, 4, 33, 14, 36], sequential probability assignment [8], learning in auctions [23, 18], robotics [16, 17], differential privacy [35], and reinforcement learning [54].
In particular, [43] presented a general framework for analyzing minimax regret against smooth adversaries in terms of a distribution-dependent sequential Rademacher complexity. Then, [36, 4] provided tight regret bounds for smoothed online learning for classification and regression respectively, under the core assumption that the base measure is known. As an important note, the notion of smoothness in terms of bounded Radon-Nikodym density with respect to the base measure can usually be generalized to general divergence balls as studied in [14].
Agnostic smoothed online learning.
Crucially, the above-mentioned works on the standard smooth online learning problem assume that the base measure $\mu$ is known to the learner. Arguably, this is a somewhat strong assumption both in practice and in theory. Knowing the base measure significantly diverges from classical results in the PAC learning setting for which knowing the distribution of the data is unnecessary, or from results from the literature on consistency which require no prior knowledge on the data-generating process. Hence, we aim to answer the following question:
Can we achieve sublinear regret for smoothed online learning without prior knowledge of the base measure? If so, which algorithm achieves the optimal excess error guarantee?
Along this direction, [15] notably showed that if the values $(y_{t})_{t\geq 1}$ are well-specified, i.e., given a function class $\mathcal{F}$, there exists some $f^{\star}\in\mathcal{F}$ such that $\mathbb{E}[y_{t}\mid x_{t}]=f^{\star}(x_{t})$ for all $t\geq 1$, then empirical risk minimization (ERM) has a regret guarantee of the form $\sigma^{-1}\sqrt{\text{comp}(\mathcal{F})\cdot T}$ for some complexity notion for the function class $\text{comp}(\mathcal{F})$ (see Theorem 3 for a complete statement). Importantly, ERM does not require any prior knowledge of the base measure. In terms of lower bounds, [4] showed that some polynomial dependency of the regret in $\sigma^{-1}$ is necessary as opposed to the setting in which $\mu$ in known for which the regret usually depends on $\ln(\sigma^{-1})$.
We focus on the general setting in which no assumptions are made on the values $(y_{t})_{t\geq 1}$ selected by the adversary, and the learner has no prior knowledge on the base measure, which we refer to as the agnostic smoothed online learning setting. As before, the goal is to achieve low regret compared to a fixed function class $\mathcal{F}$.
Our contributions.
We answer positively to the previous question by providing a proper algorithm R-Cover (Recursive Covering) based on recursive coverings that achieves the optimal regret guarantee in classification for function classes $\mathcal{F}$ with finite VC dimension up to logarithmic factors (Theorem 4), and sublinear regret in regression for function classes with standard fat-shattering dimension growth (Theorem 6). To the best of our knowledge, this is the first algorithm with sublinear regret guarantees for the general agnostic online learning problem without prior knowledge of the base measure. We note that R-Cover also does not require the knowledge of the smoothness parameter $\sigma$.
Our main result is easiest to present for classification. Namely, when $\mathcal{F}:\mathcal{X}\to\{0,1\}$ has VC dimension $d$, we prove that R-Cover achieves the following regret guarantee:
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right]=\tilde{\mathcal{O}}\left(\sqrt{\frac{dT}{\sigma}}\right),$
where $\tilde{\mathcal{O}}$ hides poly-logarithmic factors in $T$ only. This matches a lower bound for VC classes up to logarithmic factors concurrently obtained by the authors from [15] (confirmed via personal communication). In particular, R-Cover has optimal dependency in $T$, $d$, but also the smoothness parameter $\sigma$. Precisely, there is a function class of VC dimension $d$ for which any learning algorithm must incur an expected regret $\sqrt{dT/\sigma}$ for some smooth adversary (Theorem 5). This lower bound holds even in the realizable setting (well-specified and noiseless) in which there exists some function $f^{\star}\in\mathcal{F}$ fixed a priori for which $y_{t}=f^{\star}(x_{t})$ for all $t\geq 1$, and the loss is fixed over time.
The proof of the regret guarantees of R-Cover crucially relies on a novel property that we prove for smooth adversaries (see Propositions 9 and 13). At the high level, this tightly bounds the possible amount of exploration of unknown regions of the instance space for smooth adversaries. This may be of broader interest for smoothed analysis without prior knowledge of the base measure, or for understanding which relaxations of the smoothness assumption could be made while preserving regret guarantees.
2 Preliminaries
2.1 Formal setup
We start by formally defining the online learning problem. Let $\mathcal{X}$ be an instance space equipped with some sigma-algebra. The function class $\mathcal{F}$ is a set of measurable functions $f:\mathcal{X}\to[0,1]$. We fix a horizon $T\geq 1$ and consider the following sequential prediction task. At each iteration $t\in[T]$,
  1. 1.
    An adversary chooses a distribution $\mu_{t}$ on $\mathcal{X}$ depending on all history, samples $x_{t}\sim\mu_{t}$ independently from the history, then chooses a $1$-Lipschitz loss function $\ell_{t}:[0,1]\to[0,1]$ depending on $x_{t}$ and the history.
  2. 2.
    The learner observes $x_{t}$ and makes a prediction $\hat{y}_{t}\in[0,1]$.
  3. 3.
    The learner observes $\ell_{t}$ and incurs the loss $\ell_{t}(\hat{y}_{t})$.
In particular, this captures the standard prediction setting in which there is a fixed $1$-Lipschitz loss $\ell:[0,1]\times[0,1]\to[0,1]$ and the loss of the learner is given as $\ell(\hat{y}_{t},y_{t})$ for some value $y_{t}$ that is revealed after the prediction $\hat{y}_{t}$. Indeed, the adversary may choose the loss $\ell_{t}(\cdot)=\ell(\cdot,y_{t})$ in step $1$. Next, we say that the learner is proper if at each iteration $t\in[T]$, before observing the query $x_{t}$, the learner first commits to a function $\hat{f}_{t}\in\mathcal{F}$ then, upon observing $x_{t}$, predicts the value $\hat{y}_{t}=\hat{f}_{t}(x_{t})$. Our proposed algorithms will enjoy this property.
The smoothness assumption constrains the choices for the distributions $\mu_{t}$ chosen by the adversary as defined below.
Definition 1 (Smooth distributions and smooth adversaries).
Let $\mu,p$ be probability measures on $\mathcal{X}$. We say that $p$ is $\sigma$-smooth with respect to $\mu$ if $\|\frac{dp}{d\mu}\|_{\infty}\leq 1/\sigma$, where $\|\cdot\|_{\infty}$ denotes the essential supremum. We say that an adversary is $\sigma$-smooth with respect to the base measure $\mu$ if for any $t\in[T]$, the distribution $\mu_{t}$ selected by the adversary in step 1 above is $\sigma$-smooth with respect to $\mu$.
The goal of the learner is to minimize their regret, that is, the excess error compared to the benchmark functions in $\mathcal{F}$. Precisely, we distinguish between the expected adaptive regret
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right],$
in which the benchmark function may depend on the specific realizations of the learning process, and the expected oblivious regret
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})\right]-\inf_{f\in\mathcal{F}}\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right],$
in which the benchmark function is fixed prior to the learning process. Adaptive benchmark are known to require significantly stronger analysis than oblivious benchmarks for smoothed online learning (e.g. see [36]).
2.2 Complexity notions for the function class and prior results
In classification, when the functions take value in $\{0,1\}$, when the instance process is i.i.d. ($\sigma=0$) it is known that in our setup, learnability is characterized by the VC dimension [49, 50, 48].
Definition 2 (VC dimension).
Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class for classification. We say that $\mathcal{F}$ shatters a set of points $\{x_{1},\ldots,x_{m}\}\subset\mathcal{X}$ if for any choice of values $\epsilon\in\{0,1\}^{m}$ there exists $f_{\epsilon}\in\mathcal{F}$ such that $f_{\epsilon}(x_{i})=\epsilon_{i}$ for all $i\in[m]$. The VC dimension of $\mathcal{F}$ is the size of the largest shattered set.
We next state the classical Sauer-Shelah’s lemma [45, 46] which bounds the size of the projection of a function class with finite VC dimension onto a set $\{x_{1},\ldots,x_{n}\}\subset\mathcal{X}$.
Lemma 1 (Sauer-Shelah’s lemma).
Let $\mathcal{F}$ be a function class from $\mathcal{X}$ to $\{0,1\}$ of VC dimension $d$. Then, for any $x_{1},\ldots,x_{n}\in\mathcal{X}$,
|$\left|\left\{(f(x_{i}))_{i\in[n]},f\in\mathcal{F}\right\}\right|\leq\sum_{i=0}^{d}\binom{n}{i}.$
In particular, the above quantity is bounded by $2n^{d}$ and if $n\geq d$ it is bounded by $\left(\frac{2en}{d}\right)^{d}$.
In the regression setting for which functions take value on the interval $[0,1]$, a scale-dependent analog characterizes the learnability of the function class $\mathcal{F}$. This is known as the fat-shattering dimension of the class [13, 37].
Definition 3 (Fat-shattering dimension).
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class for regression. Fix a scale $\alpha>0$. We say that $\mathcal{F}$ $\alpha$-shatters a set of points $\{x_{1},\ldots,x_{m}\}\subset\mathcal{X}$ if there exist values $s_{1},\ldots,s_{m}\in[0,1]$ such that for any choice of signs $\epsilon\in\{\pm 1\}^{m}$ there exists $f_{\epsilon}\in\mathcal{F}$ such that $\epsilon(f_{\epsilon}(x_{i})-s_{i})\geq\alpha$ for all $i\in[m]$. The fat-shattering dimension of $\mathcal{F}$ at scale $\alpha>0$, denoted $\mathsf{fat}_{\mathcal{F}}(\alpha)$ is the size of the largest $\alpha$-shattered set.
Generalizing Sauer-Shelah’s lemma, it is known that the fat-shattering dimension can be used to bound the size of empirical covers regression function classes. Before stating the bound, we formally define the notion of covering set and covering numbers. We voluntarily restrict these notions to the empirical infinite norm, which is sufficient for this work.
Definition 4 (Covering set and covering numbers).
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class for regression. Fix a set $S=\{x_{1},\ldots,x_{n}\}\subset\mathcal{X}$ and $\epsilon\geq 0$. We say that $\mathcal{C}\subset\mathcal{F}$ is an $\epsilon$-cover of $\mathcal{F}$ on $S$ if for all $f\in\mathcal{F}$ there exists $g\in\mathcal{C}$ such that
$\max_{i\in[n]}|f(x_{i})-g(x_{i})|\leq\epsilon.$
The $\epsilon$-covering number of $\mathcal{F}$ on $S$, denoted $\mathcal{N}(\mathcal{F};\epsilon,S)$ is the size of the smallest $\epsilon$-cover of $\mathcal{F}$ on $S$.
The following result bounds these covering numbers similarly to Sauer-Shelah’s Lemma 1.
Theorem 2 (Theorem 4.4 from [44]).
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class and let $S\subset\mathcal{X}$ be a finite set. Then, for any $\alpha\in(0,1)$ there are constants $c,C>0$ such that
$\ln\mathcal{N}(\mathcal{F};\epsilon,S)\lesssim\mathsf{fat}_{\mathcal{F}}(c\alpha\epsilon)\ln^{1+\alpha}\left(\frac{C|S|}{\mathsf{fat}_{\mathcal{F}}(c\epsilon)\epsilon}\right).$
To state some of our results, we also need to define the Wills functional [53, 30] of $\mathcal{F}$, which was first introduced in the context of lattice point enumeration. The definition below uses the formulation from [15].
Definition 5 (Wills functional).
Fix values $Z_{1},\ldots,Z_{m}\in\mathcal{X}$ and let $\xi=(\xi_{1},\ldots,\xi_{m})$ be a vector of i.i.d. standard Gaussian random variables. The Wills functional of $\mathcal{F}$ on $Z_{1},\ldots,Z_{m}$ is defined as
$W_{m,Z}(\mathcal{F}):=\mathbb{E}_{\xi}\left[\exp\left(\sup_{f\in\mathcal{F}}\sum_{i=1}^{m}\xi_{i}f(Z_{i})-\frac{1}{2}f(Z_{i})^{2}\right)\right].$
Note that the above definition depends on the choice of $Z_{1},\ldots,Z_{m}$. For simplicity we may omit this dependency—most of the time we will take its expectation for $Z_{1},\ldots,Z_{m}\overset{iid}{\sim}\mu$. The properties of the Wills functional have been extensively studied [53, 30, 39, 41]. We refer to [41] for detailed connections with metric complexities and universal coding. We give in Appendix A a brief overview of links between Wills functional and other more standard measures complexities that are most relevant to this work, including the VC and fat-shattering dimensions.
In particular, we have $\ln W_{m}(\mathcal{F})\lesssim d\ln m$ for classes $\mathcal{F}$ with finite VC dimension. [41] also showed that we can bound $\ln W_{m}(\mathcal{F})\leq\mathcal{G}_{m}(\mathcal{F})$ where $\mathcal{G}_{m}(\mathcal{F})$ is the Gaussian complexity of $\mathcal{F}$ (see Appendix A for a definition). Last, [15] showed that having $\ln W_{m}(\mathcal{F})=o(m)$ is necessary and sufficient to ensure learnability with polynomially many samples when the data is i.i.d.
Now that we have defined the Wills functional, we can formally state the main result from [15] which shows that empirical risk minimization (ERM) achieves sublinear regret without knowledge of the base measure for the specific well-specified setting.
Theorem 3 (Theorem 1 of [15]).
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function a function class. Consider the squared loss regression setting in which $\ell_{t}(\cdot)=(\cdot-y_{t})^{2}$ for a value $y_{t}\in\mathbb{R}$. Suppose that there exists some function $f^{\star}\in\mathcal{F}$ such that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ and that the values are given via $y_{t}=f^{\star}(x_{t})+\eta_{t}$ where $\eta_{t}\mid\mathcal{H}_{t-1}$ is a mean-zero subgaussian random variable with variance proxy $\nu^{2}$. Then, ERM makes predictions $\hat{y}_{t}$ such that
$\mathbb{E}\left[\sum_{t=1}^{T}(\hat{y}_{t}-f^{\star}(x_{t}))^{2}\right]\leq\frac{20\ln^{3}T}{\sigma}\sqrt{T(1+\nu)(1+\ln\mathbb{E}_{\mu}\left[W_{2T\ln(T)/\sigma}(256\mathcal{F})\right])}.$
2.3 Further definitions and notations
We define the notion of tangent sequence [24] which will be useful within the proofs.
Definition 6 (Tangent sequence).
Let $(Z_{t})_{t\geq 1}$ be a sequence of random variables adapted to a filtration $(\mathcal{F}_{t})_{t\geq 1}$. A tangent sequence $(Z_{t}^{\prime})_{t\geq 1}$ is a sequence of random variables such that $Z_{t}$ and $Z_{t}^{\prime}$ are i.i.d. conditionally on $\mathcal{F}_{t-1}$.
Throughout this work, we will use this notation with primes to denote tangent sequences. We also denote by $\mathcal{H}_{t}$ the history at the end of iteration $t\geq 0$ of the learning process, which is the sigma-algebra generated by $(x_{l},\hat{y}_{l},\ell_{l})_{l\leq t}$. In particular, $x_{t}\mid\mathcal{H}_{t-1}\sim\mu_{t}$ where $\mu_{t}$ is the distribution selected by the adversary in step 1 of the learning process. We use the notation $[T]:=\{1,\ldots,T\}$. We write $\lesssim$ to signify that the inequality holds up to universal constants. Last unless mentioned otherwise, the notation $\tilde{\mathcal{O}}$ only hides poly-logarithmic factors in $T$.
3 Main results
We start by giving our main regret guarantees for our algorithm R-Cover in Section 3.1. We then construct in Section 3.2 the algorithm R-Cover instantiated for classification in which case $\mathcal{F}$ is a function class with finite VC dimension $d$. The classification case already provides most of the necessary intuitions and for ease of presentation, we defer the construction of the algorithm in the general regression case to Section 5.1. Last, R-Cover requires a specific variant for a learning with expert advice algorithm which is defined in Section 3.3.
3.1 Main regret guarantees
While our analysis provides regret bounds for general regression function classes, these are more easily stated for classification. In particular, we obtain the following result.
Theorem 4.
Fix $T\geq 1$. Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class with VC dimension $d$. Suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. Then, R-Cover makes predictions $\hat{y}_{t}$ such that
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right]\leq C\ln^{5/2}T\sqrt{\frac{dT}{\sigma}},$
for some universal constant $C>0$.
As a by-product of the analysis, we also provide a high-probability version of the above expected adaptive regret bound (see Eq. 30). Note that compared to the regret bound Theorem 3 which becomes $\sigma^{-1}\ln^{7/2}(T)\sqrt{dT}$ for VC classes, our regret bound holds for adversarial values $(y_{t})_{t\in[T]}$ and has an improved dependency in $\sigma$: it grows as $1/\sqrt{\sigma}$ instead of $1/\sigma$. In particular, this yields non-trivial regret bounds for any $\sigma\in[d/T,1]$. Our regret bound for R-Cover is complemented by a matching lower bound up to logarithmic factors, which holds even in the realizable noiseless setting. Confirmed by personal communication, the authors from [15] generalized their lower bound for the regret empirical risk minimization (ERM) (Theorem 3) to general algorithms for VC classes, leading to the same result as below. We include the proof in Appendix C for completeness. The proof strategy is also of independent interest and can be used to show that some of the properties we develop on smooth adversaries (Propositions 9 and 13) are essentially tight. We refer to Section 4.1 for further discussion.
Theorem 5.
Fix $d\geq 1$. There exists a function class $\mathcal{F}:\mathcal{X}\to\{0,1\}$ with VC dimension $d$ such that for any $\sigma\in(0,1)$, $T\geq 1$, and any learning algorithm, there is a function $f^{\star}\in\mathcal{F}$ and a $\sigma$-smooth adversary such that the responses are realizable, that is, $y_{t}=f^{\star}(x_{t})$ for all $t\in[T]$, and denoting by $\hat{y}_{t}$ the predictions of the algorithm,
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}[\hat{y}_{t}\neq f^{\star}(x_{t})]\right]\geq\min\left(\frac{1}{12}\sqrt{\frac{dT(1-\sigma)}{\sigma}},\;\frac{T}{24}\right).$
As a remark, R-Cover uses a somewhat complex recursive construction to achieve the optimal regret guarantee from Theorem 4. Achieving (worse) sublinear regret without prior knowledge of the base measure with a simpler algorithm is nevertheless possible. In Section 4.1 we describe a very simple and intuitive Cover algorithm which essentially corresponds to the single-depth version of R-Cover and for instance enjoys a $\approx T^{2/3}$ regret guarantee in classification. We refer to Section 4.1 for details on this result which may be of independent interest. Obtaining regret guarantees in regression for Cover is also possible with the same tools developed for R-Cover and we omit details for simplicity.
We next turn to the general regression setting. At the high level, our algorithm for classification is generalized to regression by constructing $\epsilon$-coverings of the function class for some scale $\epsilon$ that is used as a parameter (for VC classes we simply use $\epsilon=0$). In practice, the optimal choice of the scale $\epsilon$ lies in $[1/T,1]$ and only depends on the growth of the fat-shattering dimensions of $\mathcal{F}$. We note, however, that tuning this parameter $\epsilon$ can be fully side-stepped by performing any learning with expert advice algorithm using as experts the algorithms R-Cover for different choice of parameters $\epsilon\in\{2^{-l},l\leq\log_{2}T\}$. The resulting algorithm would enjoy the same regret guarantees as for the optimally-tuned algorithm.
The full version of our regret bound is stated in Theorem 14. For readability, we instantiate the bound for standard growth scenarios for the fat-shattering dimension of $\mathcal{F}$.
Theorem 6.
Fix $T\geq 1$. Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class. and suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. There exists a universal constant $C>0$ such that we have the following bounds on the oblivious regret of R-Cover, where we denote by $\hat{y}_{t}$ the predictions of the algorithm.
If $\mathsf{fat}_{\mathcal{F}}(r)\leq d\ln\frac{1}{r}$ for all $r>0$, then R-Cover run with the parameter $\epsilon=1/T$ yields
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})\right]-\inf_{f\in\mathcal{F}}\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right]\leq C\ln^{3}T\sqrt{\frac{dT}{\sigma}}.$
If $\mathsf{fat}_{\mathcal{F}}(r)\lesssim r^{-p}$ for $p>0$, then R-Cover run with the parameter $\epsilon=\left(\frac{\ln T}{T}\right)^{\frac{1}{p+1}}$ yields
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})\right]-\inf_{f\in\mathcal{F}}\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right]\leq C\frac{\ln^{3}T}{\sqrt{\sigma}}\cdot T^{1-\frac{1}{2(p+1)}}\left(1+\tilde{O}(\sigma^{-\frac{1}{2}}T^{-\frac{\min(p,1)}{2(p+1)(p+2)}})\right),$
where $\tilde{O}$ only hides logarithmic factors in $T$.
As for Theorem 4, our analysis also provides high-probability versions of the bounds in Theorem 6 (see Theorem 14). Note that the guarantee for classification from Theorem 4 bounds the expected adaptive regret, while in the regression case, Theorem 6 bounds the expected oblivious regret. We leave open the question of whether one can achieve guarantees for the adaptive regret in this case.
3.2 Recursive construction of R-Cover for classification
In its simplest form, R-Cover subdivides the horizon $[T]$ into $K$ equal-length epochs and uses a learning with expert advice algorithm on each epoch on a subset of functions from $\mathcal{F}$ that are representative from the data observed in previous epochs. For this simpler version, we can for instance use the classical Hedge algorithm [19] on the projection of $\mathcal{F}$ on the queries observed on previous epochs. While we can show that this algorithm already achieves a sublinear regret (see Section 4.1 and Theorem 8 for a detailed discussion), to achieve a $\approx\sqrt{T}$ regret, we need to use a recursive construction, which we parameterize by a depth parameter $P\geq 0$. Intuitively, the approach mentioned above corresponds to the depth-$1$ algorithm.
[…]
Recursive construction.
1
 Input: depth $P\geq 0$, start and end times $T_{0}\leq T_{1}$ satisfying $T_{1}-T_{0}\geq 2^{P}$, realizable labeled dataset $S\subset\mathcal{X}\times\{0,1\}$
2
3
4if $P=0$ then
5       Fix $f_{S}\in\mathcal{F}$ realizing dataset $S$ and predict $\hat{y}_{t}=f_{S}(x_{t})$ for all $t\in(T_{0},T_{1}]$
6else
7       Fix $f_{S}\in\mathcal{F}$ realizing dataset $S$ and let $T_{1/2}:=\left\lfloor\frac{T_{0}+T_{1}}{2}\right\rfloor$
8      for $\alpha\in\{0,1/2\}$ (epoch $(T_{\alpha},T_{\alpha+1/2}])$ do
9             After iteration $T_{\alpha}$, construct all distinct realizable datasets $S_{1},\ldots,S_{r}\subset\mathcal{X}\times\{0,1\}$ obtained by adding labeled points $(x_{t},y_{t})_{t\in(T_{0},T_{\alpha}]}$ for queries from previous epoch to previous dataset $S$
[…]
11       end for
12
13 end if
14
[…]
For the base depth $P=0$, given start and end times $T_{0}<T_{1}$ and a dataset $S$, the algorithm simply selects any arbitrary function $f_{S}\in\mathcal{F}$ that agrees on the query set, that is $f_{S}(x)=y$ for all $(x,y)\in S$, and uses it as prediction at all times in $[T]$.
[…]
As an important remark, because $\mathcal{F}$ has VC dimension $d$, we always have $r\leq 2T^{d}+1$ from Sauer-Shelah’s Lemma 1. The additional expert comes from the fact that we also added $f_{S}$ as expert.
Final algorithm.
[…]
3.3 Learning with expert advice algorithm
Instead of using the standard exponentially weighted algorithm for learning with expert advice, we use a specific variant. We briefly recall the setup for prediction with $K$ experts and fixed horizon $T$ that is relevant for our present discussion. At each iteration $t\in[T]$, the environment chooses losses $\ell_{t,i}$ for each experts $i\in[K]$. The learner then selects an expert $\hat{i}_{t}\in[K]$ potentially randomly without knowledge of the losses at time $t$. Last, all losses at time $t$ are revealed to the learner and they incur the loss $\ell_{t,\hat{i}_{t}}$ from the selected expert. The goal of the learner is to minimize its regret compared to the performance of any fixed expert:
$Reg(T):=\sum_{t=1}^{T}\ell_{t,\hat{i}_{t}}-\min_{i\in[K]}\sum_{t=1}^{T}\ell_{t,i}.$
The classical exponentially weighted forecaster or Hedge algorithm (see e.g. [19]) with parameter $\eta>0$ proceeds as follows. At time $t$, it computes the cumulative regret compared to each expert up to time $t$: $R_{t-1,i}:=\sum_{l=1}^{t-1}\ell_{l,\hat{i}_{l}}-\ell_{l,i}$ for all $i\in[K]$. It then randomly samples $\hat{i}_{t}\sim p_{t}$ where the distribution $p_{t}=(p_{t,i})_{i\in[K]}$ is defined via exponential weights
$p_{t,i}:=\frac{e^{\eta R_{t-1,i}}}{\sum_{j\in[K]}e^{\eta R_{t-1,j}}}.$
We next denote by $\mathcal{F}_{t}=(\ell_{l,i},l\leq t,i\in[K],\hat{i}_{l},l<t)$ the history up to time $t$ included. The exponentially weighted forecaster with learning parameter $\eta$ enjoys the following classical bound (see e.g. [19, Corollary 2.2]):
$PReg(T):=\sum_{t=1}^{T}\mathbb{E}_{\hat{i}_{t}}[\ell_{t,\hat{i}_{t}}\mid\mathcal{F}_{t}]-\min_{i\in[K]}\sum_{t=1}^{T}\ell_{t,i}\leq\frac{\ln K}{\eta}+\frac{T\eta}{2}.$
We will refer to the quantity on the left-hand side as the pseudo-regret $PReg(T)$. Using the standard choice of parameter $\eta=\sqrt{2\ln K/T}$, and assuming that the losses all have values in $[0,1]$, the previous equation directly gives an expected bound on the regret $\mathbb{E}[Reg(T)]\lesssim\sqrt{T\ln K}$. However, for our purposes, we need a refinement of this bound. Using [19, Theorem 2.1], we can derive the following bound
$PReg(T)\leq\frac{\ln K}{\eta}+\frac{\eta}{2}\sum_{t=1}^{T}\sum_{i\in[K]}p_{t,i}r_{t,i}^{2},$ (1)
where $r_{t,i}:=\ell_{t,\hat{i}_{t}}-\ell_{t,i}$ is the instantaneous regret of the forecaster compared to expert $i$ at time $t$. For convenience, let us denote
$\Delta_{T}:=\sum_{t=1}^{T}\sum_{i\in[K]}p_{t,i}r_{t,i}^{2}=\sum_{t=1}^{T}\mathbb{E}_{\hat{i}_{t}}[r_{t,\hat{i}_{t}}^{2}\mid\mathcal{F}_{t}].$
Eq. 1 yields a tighter bound than the standard regret bound if one selects $\eta\approx\sqrt{\ln K/\Delta_{T}}$ instead of the standard choice $\eta\approx\sqrt{\ln K/T}$. Achieving the corresponding bound without a prior knowledge of $\Delta_{T}$ can be easily performed via the standard doubling trick. Precisely, we use the exponentially weighted forecaster with initial parameter $\eta_{1}\approx\sqrt{2\ln K}$ until $\Delta_{t}\geq 1$, then restart the algorithm with a parameter $\eta_{2}\approx\eta_{1}/2$ until $\Delta_{t}\geq 4$. We continue the process by always restarting the algorithm with a quadrupled threshold for $\Delta$ and a corresponding parameter $\eta>0$ (roughly halved). The precise algorithm is given in Algorithm 2, which is the exponentially weighted forecaster variant that we use for our algorithm R-Cover. This variant enjoys the following pseudo-regret bound, whose proof is given in Appendix B.
1
 Input: number of experts $K$
2
3 Let $k=1$, $\Delta_{max,1}=1$, $\eta_{1}=\sqrt{2\ln K/(\Delta_{max,1}+1)}$
4Initialize $R_{0,i}=0$ for all $i\in[K]$, and $\Delta_{1}=0$
5for $t\geq 1$ do
6       Let $p_{t,i}=\frac{e^{\eta_{k}R_{t-1,i}}}{\sum_{j\in[K]}e^{\eta_{k}R_{t-1,j}}}$ and sample $\hat{i}_{t}\sim p_{t}$ independently from history
7      Observe $\ell_{t,i}$ for $i\in[K]$, let $r_{t,i}=\ell_{t,\hat{i}_{t}}-\ell_{t,i}$ and $R_{t,i}=R_{t-1,i}+r_{t,i}$ for $i\in[K]$
8      Update $\Delta_{k}\leftarrow\Delta_{k}+\sum_{i\in[K]}p_{t,i}r_{t,i}^{2}$
9      if $\Delta_{k}>\Delta_{max,k}$ then
10             Set $\Delta_{max,k+1}=4\Delta_{max,k}$ and $\eta_{k+1}=\sqrt{2\ln K/(\Delta_{max,k+1}+1)}$
11            Reset $R_{t,i}=0$ for all $i\in[K]$, $\Delta_{k+1}=0$, and $k\leftarrow k+1$
12       end if
13
14 end for
15
Algorithm 2 Adaptive exponentially weighted forecaster A-Exp
Lemma 7.
Suppose that all losses lie in $[0,1]$. Then, the pseudo-regret of the adaptive exponentially weighted forecaster A-Exp satisfies
$PReg(T)\leq 8\sqrt{\max(\Delta_{T},1)\ln K},\quad T\geq 1.$
Further, for $T\geq 1$ and $\delta\in(0,1)$, with probability at least $1-\delta$ we have
$Reg(T)\leq 12\sqrt{\max(\Delta_{T},1)\ln K}+2\ln\frac{1}{\delta}.$
4 Technical overview
As discussed above, the classification setting will be mostly sufficient to present our main proof ideas. Hence, in this section we mostly focus on this case, that is, we suppose that $\mathcal{F}:\mathcal{X}\to\{0,1\}$ has VC dimension $d$.
4.1 A simple algorithm for a weaker regret guarantee
To motivate the form of R-Cover, we first consider a significantly simpler algorithm which essentially corresponds to R-Cover with depth $1$. In this simplest form, R-Cover subdivides the horizon $[T]$ into $K$ equal-length epochs and a learning with expert advice algorithm on each epoch on the projection of the function class $\mathcal{F}$ on query points $x_{t}$ from prior epochs. For instance, we can use the classical exponentially weighted forecaster algorithm (e.g. see [19]). This simplified algorithm which we call Cover is summarized in Algorithm 3.
1
 Input: horizon $T$, number of epochs $K\leq T$
2
3
4Let $T_{k}=\left\lfloor k\frac{T}{K}\right\rfloor$ for $k\in\{0,\ldots,K\}$.
5for $k\in[K]$ do
6       Construct a minimal-size cover $S_{k}\subset\mathcal{F}$ such that for any $f\in\mathcal{F}$ there exists $g\in S_{k}$ with $f(x_{s})=g(x_{s})$ for $s\in[T_{k-1}]$
7      For iterations $t\in(T_{k-1},T_{k}]$, run any learning with expert advice algorithm (e.g. Hedge) with expert set $S_{k}$
8 end for
9
Algorithm 3 Construction of the Cover algorithm
We can show that with a convenient choice of the number of epochs $K\approx T^{1/3}$, Cover already achieves a $\approx T^{2/3}$ regret guarantee without any prior knowledge on the distribution $\mu$. Given the simplicity of Cover, this result may be of independent interest.
Theorem 8.
Fix $T\geq 1$. Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class with VC dimension $d$. Suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. Then, Cover run with parameter $K=\lfloor\ln T\cdot(T/d)^{1/3}\sigma^{-2/3}\rfloor$ makes predictions $\hat{y}_{t}$ such that
$\mathbb{E}\left[\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\right]\leq C\ln^{2}T\left(\frac{dT^{2}}{\sigma}\right)^{1/3}.$
for some universal constant $C>0$.
We formally prove this result in Appendix D. In this section, our goal is mostly to give key intuitions about the underlying strategy for the full algorithm R-Cover. To give some insights into why Cover already achieves sublinear regret, note that if the queries prior to some epoch $(T_{k-1},T_{k}]$ are “representative” of the queries during this epoch, then the cover $S_{k}$ constructed at the beginning of the epoch (line 3 of Algorithm 3) is a good representative set of relevant functions. Naturally, this holds if the underlying process $(x_{t})_{t\in[T]}$ is i.i.d.—that is $\sigma=0$. The crux of our analysis is to show that when the adversary is $\sigma$-smooth this still holds in an amortized sense. Note that it is not true that the queries $(x_{t})_{t\leq T_{k-1}}$ observed prior to some epoch $(T_{k-1},T_{k}]$ are always representative of the queries during that epoch. Indeed, a $\sigma$-smooth adversary can for instance decide to have the sequence of distributions $(\mu_{t})_{t\in[T]}$ adaptively switch from one distribution to a completely unrelated one up to $\left\lfloor 1/\sigma\right\rfloor$ times. However, we show that the number of epochs for which prior queries $(x_{t})_{t\leq T_{k-1}}$ are not representative of the queries on the epoch $(T_{k-1},T_{k}]$ is bounded.
To quantify the notion of “representativeness”, we introduce the following quantity, which essentially quantifies the maximum $\ell_{1}$ discrepancy between queries observed until some time $t_{0}<t$ and the query made at time $t$ on the function class $\mathcal{F}$. For any $0\leq t_{0}<t\leq T$, we define
$\gamma_{t_{0}}(t):=\sup_{\begin{subarray}{c}f,g\in\mathcal{F}\text{ s.t.}\\ f(x_{s})=g(x_{s}),\,s\in[t_{0}]\end{subarray}}\mathbb{P}\left(f(x_{t})\neq g(x_{t})\mid\mathcal{H}_{t-1}\right)=\sup_{\begin{subarray}{c}f,g\in\mathcal{F}\text{ s.t.}\\ f(x_{s})=g(x_{s}),\,s\in[t_{0}]\end{subarray}}\mathbb{P}_{x\sim\mu_{t}}\left(f(x)\neq g(x)\right),$ (2)
where we recall that $\mathcal{H}_{t-1}$ denotes all history available until the end of iteration $t-1$. Intuitively, if the queries prior to $t_{0}$ were representative of the query at time $t$, then the empirical projection of $\mathcal{F}$ onto the query set $(x_{t})_{t\leq t_{0}}$ should reasonably cover $x_{t}\sim\mu_{t}$ and as a result $\gamma_{t_{0}}(t)$ would be smaller.
One of our main contributions for the analysis of smoothed adversaries is the following result which bounds the number of epochs on which prior history is not representative. How the epochs are constructed is very flexible: we used a fixed schedule for Cover and R-Cover but randomized epochs are also possible, which may be useful for improved regret bounds in the regression case. The proof uses some key results from [15].
Proposition 9.
Let $T\geq 2$ and $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class with VC dimension $d$.
Consider any online mechanism to construct epochs $(T_{k-1},T_{k}]$ for $k\in[K]$. That is, let $(T_{k})_{k\geq 0}$ be random times such that (1) $T_{0}=0$, (2) for all $k\geq 1$, $T_{k}\mid\{T_{k-1},T_{k-1}<T\}$ is a stopping time adapted to the filtration $(\mathcal{H}_{t})_{t\geq T_{k-1}}$, and (3) for all $k\geq 1$ almost surely, $T_{k-1}<T_{k}\leq T$ conditionally on $T_{k-1}<T$. Let $K\leq T$ denote the first index such that $T_{K}=T$.
Fix any parameters $q,\delta\in(0,1]$ and denote $w(T,\delta):=d\ln\left(\frac{T}{\sigma}\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}+2$. Then, with probability at least $1-\delta$,
|$\left|\left\{k\in[K]:\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)\cdot\mathbbm{1}[\gamma_{T_{k-1}}(t)\geq q]\geq w(T,\delta)\right\}\right|\leq C\frac{\ln^{2}T}{q\sigma},$
for some universal constant $C\geq 1$. For a bound in expectation we can simply take $w(T):=d\ln\frac{T}{\sigma}+2$.
Proposition 9 shows that up $\tilde{\mathcal{O}}(1/(q\sigma))$ epochs, we only pay at most a price $w(T,\delta)$ during each epoch $(T_{k-1},T_{k}]$ for the times $t\in(T_{k-1},T_{k}]$ when the cover constructed from queries prior to this epoch was not representative of query $x_{t}$ by some threshold $q$. Here, we largely view $w(T,\delta)$ as a reasonable price to pay on each epoch. Hence, intuitively, we can consider that up to $\tilde{\mathcal{O}}(1/(q\sigma))$ epochs, the cover constructed from queries on prior epochs is always representative from the queries on the epoch up to threshold $q$.
We emphasize that up to the logarithmic factors, Proposition 9 is tight in the following sense. For any choice of the online mechanism to construct epochs and threshold $q$, a $\sigma$-smooth adversary can ensure that for $\mathcal{O}(1/(q\sigma))$ epochs $(T_{k-1},T_{k}]$, queries on prior epochs are not representative up to threshold $q$ from all times in $(T_{k-1},T_{k}]$. We detail below the scenarios for which Proposition 9 is tight. We believe that these essentially captures all possible attack behaviors of a smooth adversary.
Because of the $\sigma$-smoothness constraint, the adversary cannot query the algorithm on completely different regions of the space $\mathcal{X}$ at each epoch. One possible strategy for the adversary, which we discussed as motivation above, is to switch distributions $\left\lfloor 1/\sigma\right\rfloor$ times during the learning process, possibly onto a completely new region of the space. This corresponds to $q=1$ in Proposition 9: at the start of $\left\lfloor 1/\sigma\right\rfloor$ epochs $(T_{k-1},T_{k}]$, the adversary switches query distributions $\mu_{t}$ and selects a distribution with support on a new region for which prior queries are irrelevant. This results in $\gamma_{T_{k-1}}(t)=1$ for all $t\in(T_{k-1},T_{k}]$.
A more refined strategy for the adversary in order to increase its number of affected epochs is to select a parameter $q$ and at the start of a new epoch $(T_{k-1},T_{k}]$, switch the query distribution as follows. They construct a new mixture distribution $\mu_{k}:=q\nu_{k}+(1-q)\mu_{0}$ where with probability $q$ the learner is queried on a new distribution $\nu_{k}$ with say completely new support compared to the history, and with probability $1-q$ the learner is queried on a base measure $\mu_{0}$ that is very similar to previous queries. This results in $\gamma_{T_{k-1}}(t)\geq q$ for all $t\in(T_{k-1},T_{k}]$. On one hand, during the epoch, the adversary could only test the learner on a fraction $q$ of “truly adversarial” queries sampled from $\mu_{k}$. On the other hand, the smoothness constraint is now easier to satisfy and we can check that the adversary can afford to corrupt $\approx 1/(q\sigma)$ epochs in this manner. This precisely corresponds to the bound from Proposition 9 up to logarithmic factors. As it turns out, this mixture strategy is in fact stronger for the adversary and choosing $q\approx 1/\sqrt{T}$ is the strategy that yields the lower bound from Theorem 5.
Remark 10.
The statement from Proposition 9 is written specifically for classification, for which analyzing the $\ell_{1}$ diameter as defined in $\gamma_{t_{0}}(t)$ in Eq. 2 is amenable. The proof of Proposition 9 requires controlling the complexity of the class $\{\mathbbm{1}[f\neq g]:f,g\in\mathcal{F}\}$ which has VC dimension bounded by $2d$ if $\mathcal{F}$ has VC dimension $d$. While the VC dimension behaves nicely with this self-difference operation, this is not the case for the fat-shattering dimension which is known to behave somewhat wildly with the addition [1].222[1] notes that the function class $\mathcal{F}$ of increasing functions on $[0,1]$ always has fat-shattering dimension one at any scale, while =$\mathcal{F}-\mathcal{F}=\{f-g:f,g\in\mathcal{F}\}$ has infinite shattering dimension at all scales. To solve this issue for the regression setting, we need to localize this difference class around an oblivious benchmark function $f^{\star}$. The localized analog of $\gamma_{t_{0}}(t)$ that we use in our proofs is defined in Eq. 16. The corresponding generalization of Proposition 9 is Lemma 13. In this general regression setting, the term in $w(T,\delta)$ from Proposition 9 depending on the VC dimension $d$ is replaced by the Wills functional of $\mathcal{F}$ which measures the complexity of the class.
With the main tool Proposition 9 at hand, we can easily prove a simpler version of Theorem 8 for the expected oblivious regret. Fix some benchmark function $f^{\star}\in\mathcal{F}$. On each epoch $k\in[K]$, Cover runs a learning with expert advice algorithm on the cover $S_{k}$, which has size $\mathcal{O}(T^{d})$ by Sauer-Shelah’s Lemma 1. Hence, using classical regret bounds (e.g. [19, Corollary 2.2]), the total expected regret incurred by these algorithms is bounded by
$C\sum_{k\in[K]}\sqrt{(T_{k}-T_{k-1})\cdot d\ln T}\lesssim\sqrt{KdT\ln T},$ (3)
for some constant $C\geq 1$, where we used Jensen’s inequality in the last inequality. Next, for each $k\in[K]$, denote by $f_{k}\in S_{k}$ the function in the cover that had the correct labeling compared to $f^{\star}$, that is:
$f_{k}(x_{t})=f^{\star}(x_{t}),\quad t\in[T_{k-1}],k\in[K].$
Because $f_{k}$ is one of the experts considered during epoch $k$, it suffices to bound the remaining regret term
$\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}(f_{k}(x_{t}))-\ell_{t}(f^{\star}(x_{t}))\leq\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\mathbbm{1}[f_{k}(x_{t})\neq f^{\star}(x_{t})].$
Taking the expectation of each term for $x_{t}$ conditionally on the history $\mathcal{H}_{t-1}$, we obtain
$\mathbb{E}\left[\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}(f_{k}(x_{t}))-\ell_{t}(f^{\star}(x_{t}))\right]\leq\mathbb{E}\left[\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)\right],$
since the functions $f_{k}$ and $f^{\star}$ agreed on all queries of previous epochs. We can then use Proposition 9 which bounds the sum for each epoch $k\in[K]$. Applying Proposition 9 for $q\geq q_{0}:=\ln^{2}(T)/(K\sigma)$ bounds the number of epochs for which this sum deviates significantly. At the high level, it implies that the quantities $\gamma_{T_{k-1}}(t)$ are roughly of order $q_{0}$ in average. After computations, we obtain
$\mathbb{E}\left[\sum_{k\in[K]}\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)\right]\lesssim\frac{\ln^{3}T}{K\sigma}\cdot T.$ (4)
Putting the two regret terms from Eqs. 3 and 4 together and optimizing over the choice of $K$ gives the same bound as Theorem 8 for the expected oblivious regret of Cover. We give some ideas about how this oblivious regret guarantee can be turned into an adaptive regret guarantee in Section 4.3.
4.2 Achieving the optimal regret using recursive covers
The main obstacle for Cover for achieving the optimal regret dependency $\sqrt{T}$ in the horizon is that it needs to balance between two competing regret terms: (1) the regret incurred by learning with expert algorithms, which usually increases with the number of epochs; and (2) the discretization error obtained by approximating the optimal function using a net constructed on prior epochs, which decreases with the number of epochs.
We use a localization strategy to increase the number of effective epochs on which a cover is recomputed. To not incur a large regret term due to the learning with expert algorithms, we introduce an adaptive variant from the classical Hedge algorithm, A-Exp, which has a regret bound depending on the some notion of difficulty of the learning with expert problem instead of a worst-case bound (see Section 3.3 for a detailed exposition). Going back to an epoch $(T_{k-1},T_{k}]$ of Cover, Proposition 9 essentially implies that during most epochs $k\in[K]$ one can bound
$\sum_{t=T_{k-1}}^{T_{k}}\gamma_{T_{k-1}}(t)\lesssim q_{0}(T_{k-1}-T_{k})$ (5)
where $q_{0}=\ln^{2}(T)/(K\sigma)$. As a result, if we restrict our search space on epoch $k$ to some functions that shared the same values on previous epoch queries $(x_{t})_{t\leq T_{k-1}}$, we expect that these would only disagree (have different predictions) for a fraction $\approx q_{0}$ of the times in epoch $k$. Using the regret guarantee from A-Exp from Lemma 7 we can then show that on epochs $k\in[K]$ for which Eq. 5 holds, performing A-Exp on a set $S\in\mathcal{F}$ of functions that agreed on previous epochs incurs a learning with expert regret
$\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}(\hat{y}_{t})-\min_{f\in S}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}(f(x_{t}))\lesssim\sqrt{q_{0}(T_{k-1}-T_{k})\ln|S|},$
with reasonable probability $1-\delta$. Here, $\hat{y}_{t}$ denotes the predictions of the learning with expert advice algorithm, and we omitted lower-order terms which may depend on the probability failure $\delta$. The regret obtained should be compared to a more classical worst-case bound of order $\sqrt{(T_{k-1}-T_{k})\ln|S|}$ for the Hedge algorithm.
This regret improvement for the regret of A-Exp leads us to the following depth-$2$ algorithm: on each epoch $(T_{k-1},T_{k}]$ we can run any learning with expert advice algorithm (say Hedge) using as experts the predictions of all Cover algorithms that are run with horizon $T_{k}-T_{k-1}$, use a fixed number of epochs, use A-Exp as expert advice algorithm (line 4 of Algorithm 3) and restrict their search space to functions in $\mathcal{F}$ that agreed on previous epoch queries $(x_{t})_{t\leq T_{k-1}}$. By Sauer-Shelah’s Lemma 1, there are at most $2T^{d}$ such experts. Optimizing the choice of number of epochs for each of the two layers yields an improved dependency in $T^{\alpha}$ for the final regret bound compared to the $1$-depth Cover algorithm in Theorem 8, for some $\alpha\in(1/2,2/3)$.
To achieve the optimal regret up to logarithmic terms, we run this strategy recursively over $\left\lfloor\log_{2}(T)/2\right\rfloor$ depths, which is R-Cover. This strategy is akin to some form of chaining at the algorithmic level. The smallest sub-epochs on the last layer have length of order $\sqrt{T}$. Note that the labeled dataset $S$ that is used as parameter in the recursive construction of R-Cover in Algorithm 1 now corresponds to the possible labelings of queries in prior epochs. In practice, the optimal depth to achieve the correct regret dependency in the adversary smoothness parameter $\sigma$ depends on $\sigma$ itself. To avoid requiring this information when implementing R-Cover, at each depth, in addition to the experts corresponding to the predictions of the next layer algorithm, we also add an expert that uses a single function as prediction ($f_{S}$ in line 7 of Algorithm 1). The rationale is that this hedges the final algorithm for all choices of depths at once. Within the proof, we may then focus on the algorithm up to a fixed $\sigma$-dependent depth.
4.3 Proof sketch for Theorem 4
Now that we have introduced the main conceptual ingredients of the proof, we give a brief sketch of the regret bound. We start by focusing on the oblivious regret compared to some fixed benchmark function $f^{\star}$. R-Cover is composed of $P$ layers. Each layer $p\in[P]$ corresponds to epochs $(T_{k-1}^{(p)},T_{k}^{(p)}]$ for $k\in[N_{p}]$. For instance, initially there is a single epoch for $p=P$ and at the last layer $p=0$ there are $\approx\sqrt{T}$ epochs. For each depth $p\in[P]$ and epoch $k\in[N_{p}]$ we consider the depth-$p$ R-Cover algorithm that was instantiated with the “correct” labeling according to $f^{\star}$. That is, we focus on the algorithm that used the dataset
$S_{k}^{(p)}=\left\{(x_{t},f^{\star}(x_{t})),t\in[T_{k-1}^{(p)}]\right\}.$
Regret decomposition.
The main point of the recursive procedure is that it allows to localize the error by focusing only on the runs of R-Cover that used these correct labeled datasets. The first step of the proof (Section 5.2) is to show that we can decompose the regret of the algorithm compared to $f^{\star}$ in the following way, where $\hat{y}_{t}$ denotes the predictions of the final algorithm. With probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))\lesssim\sum_{p=p_{0}}^{P}\sum_{k\in[N_{p}]}\underbrace{\sqrt{\max\left(\Delta_{k}^{(p)},2\right)d\ln T}}_{Reg_{k}^{(p)}}\\ +\sum_{k\in[N_{p_{0}}]}\underbrace{\sum_{t=T_{k-1}^{(p_{0})}}^{T_{k}^{(p_{0})}}\ell_{t}\left(f_{k,S}^{(p_{0})}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))}_{\Lambda_{k}^{(p_{0})}}+N_{p_{0}}\ln\frac{1}{\delta}.$ (6)
The first term of Eq. 6 corresponds to the regret accumulated along the localization trajectory for running the learning with expert advice algorithm A-Exp. Up to minor details, here $Reg_{k}^{(p)}$ corresponds to the bound on the regret incurred by A-Exp for the depth-$p$ algorithm run during epoch $k\in[N_{p}]$, which is guaranteed by Lemma 7. The quantity $\Delta_{k}^{(p)}$ is the same as that which appears in Lemma 7 and measures the difficulty of the learning with expert problem on epoch $k$ at depth $p$. Technically, the bound from Lemma 7 also includes a failure probability term which accumulated over the complete trajectory corresponds to the term $N_{p_{0}}\ln\frac{1}{\delta}$. This can be viewed as a lower order term.
The second term of Eq. 6 intuitively corresponds to the excess error of a learner that at the beginning of each depth-$p_{0}$ epoch $k\in[N_{p_{0}}]$ has access to an oracle which reveals the values of the optimal function $f^{\star}$ on all prior epoch queries $(x_{t})_{t\leq T_{k-1}^{(p_{0})}}$. Here, we use the notation $f_{k,S}^{(p_{0})}$ to denote the base function $f_{S}$ that was constructed either line 2 or 4 of Algorithm 1 during the run of the depth-$p_{0}$ algorithm R-Cover on epoch $k$ using the correct labeling $S_{k}^{(p_{0})}$. The quantity $\Lambda_{k}^{(p_{0})}$ corresponds to the excess error of this base function $f_{k,S}^{(p_{0})}$ compared to $f^{\star}$ on the epoch $k$.
Bounding each regret term via smoothness.
The next step of the proof is to bound each regret term and more precisely to bound the terms $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$. Using the same arguments as described in Section 4.1 when bounding the excess error of Cover on each epoch, we can show that the quantity
$\Gamma_{k}^{(p)}:=\sum_{t=T_{k-1}^{(p)}+1}^{T_{k}^{(p)}}\gamma_{T_{k-1}^{(p)}}(t)$
bounds both $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$ up to constant factors. Within the general proof for regression, these need to be bounded by different quantities $\Gamma_{k}^{(p,2)}$ and $\Gamma_{k}^{(p,1)}$ respectively in which $\gamma(t)$ is replaced by the $\ell_{2}$ and $\ell_{1}$ deviations from $f^{\star}$ (see Lemma 12 for the precise bound). In classification, because values lie in $\{0,1\}$ both norms are identical hence we can use a single quantity $\Gamma_{k}^{(p)}$.
The terms $\Gamma_{k}^{(p)}$ can now be bounded using Proposition 9, which is the only step so far that required the smoothness assumption on the adversary. For regression, the generalization of this result is Lemma 13. The proof of this main lemma is given in Section 5.5. In this full form, the guarantee obtained on $\Gamma_{k}^{(p,1)}$ and $\Gamma_{k}^{(p,2)}$ depends on the scale $\epsilon$ of the cover constructed at the beginning of each epoch (in classification we used $\epsilon=0$ in Algorithm 1). Naturally, the guarantee degrades as $\epsilon$ grows.
Putting everything together gives a high-probability bound on the regret of the algorithm compared to $f^{\star}$ of the same order as the desired adaptive bound from Theorem 4 (see Proposition 15). The corresponding bound in the general regression case is Theorem 14.
From oblivious to adaptive regret guarantees
The last step of the proof of Theorem 4 is to obtain adaptive regret bounds from high-probability oblivious regret bounds. This uses tools from prior works on smoothed online learning, and in particular [36].
We first construct a cover of the function class $\mathcal{F}$ with respect to the base measure $\mu$ and aim to have low regret compared to functions in this cover. Precisely, we construct a subset of $\mathcal{F}$ such that for all $f\in\mathcal{F}$ there exists $h\in\mathcal{H}$ with
$\mathbb{P}_{x\sim\mu}(f(x)\neq h(x))\leq\epsilon.$
Since $\mathcal{F}$ has VC dimension $d$, we can ensure $\ln|\mathcal{H}|\leq 2d\ln(e^{2}/\epsilon)$ (see [32] or [12, Lemma 13.6]). Using the union bound over the high-probability oblivious regret bounds from the previous section, we can ensure that R-Cover has low regret compared to all functions within the cover. We note that contrary to [36], this covering construction is only for proof purposes and is not performed within the algorithm R-Cover. In fact, since $\mu$ is unknown in our setting, computing such a cover is impossible, as exemplified by the lower bound Theorem 5.
It then only remains to show that $\mathcal{H}$ is also a good cover on the queries made by the smooth adversary $(x_{t})_{t\in[T]}$. Note that this would be immediate if these queries are i.i.d. sampled from $\mu$ from standard VC uniform convergence bounds. To reduce to the i.i.d. case, [36] show the following coupling lemma.
Lemma 11 ([36, 4]).
Let $(X_{t})_{t\in[T]}$ be $\sigma$-smooth with respect to $\mu$. Then for all $k\geq 1$, there exists a coupling of $(X_{t})_{t\in[T]}$ with random variables $\{Z_{t,j},t\in[T],j\in[k]\}$ such that the $Z_{t,j}\overset{iid}{\sim}\mu$ and on an event $\mathcal{E}_{k}$ of probability at least $1-Te^{-\sigma k}$, we have $X_{t}\in\{Z_{t,j},j\in[k]\}$ for all $t\in[T]$.
In particular, it suffices for the cover to perform well on all queries $Z_{t,j}$ for $t\in[T]$ and $j\in[k]$ for some $k\approx\sigma^{-1}\ln T$, which are i.i.d. This gives the desired adaptive regret bound by using VC uniform convergence bounds on the i.i.d. variables $Z_{t,j}$.
5 Regret analysis
In this section, we first describe our full algorithm R-Cover for regression in Section 5.1 then prove our main results Theorems 4 and 6 in the rest of the section.
5.1 General recursive procedure for regression
In this section, we generalize the algorithm given for classification in Algorithm 1 to handle general regression function classes $\mathcal{F}$. Note that at the beginning of each new epoch, R-Cover effectively computes a 00-cover of the previously observed dataset. In the regression setting, we instead compute an $\epsilon$-cover for some $\epsilon>0$ of the functions within the class $\mathcal{F}$ centered around a reference function $f_{0}\in\mathcal{F}$. This effectively restricts the search space of the algorithm to the neighborhood of $f_{0}$, and replaces the labeled dataset $S$ from Algorithm 1 in the classification case (these will be equivalent in this case). Instead of using a single reference function $f_{0}$, we use a sequence of reference functions which will correspond to reference functions from previous depths. This is used to ensure that the search space within $\mathcal{F}$ of sub-algorithms (akin to line 7 of Algorithm 1) are consistent with the search space of algorithm calls from previous depths. These reference functions $f_{i}:\mathcal{X}\to[0,1]$ are stored together with the start time of their corresponding algorithm call $t_{i}\in[T]$, within a set $S$.
[…]
$B_{f_{0}}(\mathcal{F};\epsilon,T_{0}):=\left\{f\in\mathcal{F}:\max_{t\in[T_{0}]}|f(x_{t})-f_{0}(x_{t})|\leq\epsilon\right\}.$ (7)
For the base depth $P=0$, the algorithm simply follows the predictions of any function within
$\mathcal{F}(S):=\bigcap_{(f_{0},t_{0})\in S}B_{f_{0}}(\mathcal{F};\epsilon,t_{0}),$ (8)
which corresponds to the search space of the algorithm. For $P\geq 1$, the algorithm defines two sub-epochs using $T_{1/2}$ exactly as in Algorithm 1. At the beginning of each epoch at time $T_{\alpha}$ for $\alpha\in\{0,1/2\}$, the algorithm constructs a minimum covering $\epsilon$-cover of the search space on the previously queried points $(x_{t})_{t\in[T_{\alpha}]}$ as defined in Definition 4. That is, we construct a set $\mathcal{C}\subset\mathcal{F}(S)$ such that for all $f\in\mathcal{F}(S)$ there exists $g\in\mathcal{C}$ such that
$\max_{t\in[T_{\alpha}]}|f(x_{t})-g(x_{t})|\leq\epsilon,$
[…]
1
 Input: depth $P\geq 0$, start and end times $T_{0}\leq T_{1}$ satisfying $T_{1}-T_{0}\geq 2^{P}$, set of reference functions $S$, scale $\epsilon$
2
3
4if $P=0$ then
5       Fix any $f_{S}\in\mathcal{F}(S)$ (see Eq. 8) and predict $\hat{y}_{t}=f_{S}(x_{t})$ for all $t\in(T_{0},T_{1}]$
6else
7       Fix any $f_{S}\in\mathcal{F}(S)$ and let $T_{1/2}:=\left\lfloor\frac{T_{0}+T_{1}}{2}\right\rfloor$
8      for $\alpha\in\{0,1/2\}$ (epoch $(T_{\alpha},T_{\alpha+1/2}])$ do
9             After iteration $T_{\alpha}$, construct a $\epsilon$-cover $\mathcal{C}$ of $\mathcal{F}(S)$ on the queries $(x_{t})_{t\in[T_{\alpha}]}$
[…]
11       end for
12
13 end if
14
[…]
$\ln\mathcal{N}(\mathcal{F}(S);\epsilon,T)\lesssim\mathsf{fat}_{\mathcal{F}}(c\alpha\epsilon)\ln^{1+\alpha}\left(\frac{CT}{\epsilon}\right).$ (9)
In the last inequality, we used the fact that $\mathsf{fat}_{\mathcal{F}(S)}(r)\leq\mathsf{fat}_{\mathcal{F}}(r)$ for all $r\geq 0$ since $\mathcal{F}(S)\subset\mathcal{F}$. Given that the covering numbers of all function classes $\mathcal{F}(S)$ are upper bounded by this quantity, in the rest of the paper, we may safely omit the dependency in $S$ to lighten the notations.
5.2 Regret decomposition
[…]
$T_{k}^{(p)}-T_{k-1}^{(p)}\in\left\{\left\lfloor\frac{T}{N_{p}}\right\rfloor,\left\lfloor\frac{T}{N_{p}}\right\rfloor+1\right\},\quad k\in[N_{p}].$ (10)
Next, we fix a function $f^{\star}\in\mathcal{F}$ that will serve as benchmark for the algorithm’s predictions. Importantly, we suppose for now that $f^{\star}$ is fixed and non-adaptive: it does not depend on the realizations of $(x_{t},y_{t})_{t\in[T]}$. We will later extend the regret bound to potentially adaptive benchmark functions $f^{\star}\in\mathcal{F}$ in the classification setting.
[…]
We start at depth $p=P$, for which there is a single epoch $k=1$. We then simply pose $f_{1}^{(P)}\in\mathcal{F}$ arbitrarily, and let $S_{1}^{(P)}:=\emptyset$, which is the reference function set used for the final algorithm. In particular we have $f^{\star}\in\mathcal{F}(S_{1}^{(P)})=\mathcal{F}$ (see the definition of $\mathcal{F}(S)$ in Eq. 8). Now suppose that we have constructed the reference functions $f_{k}^{(p)}$ and the set $S_{k}^{(p)}$ for some $p\in[P]$ and all $k\in[N_{p}]$ such that
$f^{\star}\in\mathcal{F}(S_{k}^{(p)}),\quad k\in[N_{p}].$
[…]
$\max_{t\in[T_{2(k-1)+l-1}^{(p-1)}]}\left|f^{\star}(x_{t})-f_{2(k-1)+l}^{(p-1)}(x_{t})\right|\leq\epsilon.$ (11)
Additionally, we construct the increased reference set
$S_{2(k-1)+l}^{(p-1)}:=S_{k}^{(p)}\cup\left\{\left(f_{2(k-1)+l}^{(p-1)},T_{2(k-1)+l-1}^{(p-1)}\right)\right\}.$
This ends the construction of the reference functions at depth $p-1$. Note that Eq. 11 exactly implies that the induction hypothesis holds at depth $p-1$.
[…]
Let us now focus on a single depth-$p$ epoch $k\in[N_{p}]$ for $p>0$. This is composed of $2$ sub-epochs $E_{2(k-1)+l}^{(p-1)}$ for $l\in[2]$. On each sub-epoch $l\in[2]$, the algorithm performs the exponentially weighted algorithm using experts that we denote $\mathcal{A}_{l,r^{\prime}}^{(p-1)}$ for $r^{\prime}\in[r_{l}^{(p-1)}]$. We also denote by $\mathcal{A}_{l,r^{\prime}}^{(p-1)}(t)$ their prediction for some time during the corresponding epoch $E^{(p-1)}_{2(k-1)+l}$. Last, we use the following notation to denote the magnitude of the expert problem at epoch $l$, where $\hat{p}_{t}$ denotes the distribution over the experts $\mathcal{A}_{l,r^{\prime}}^{(p-1)}(t)$ that was used by A-Exp at time $t$:
$\tilde{\Delta}_{k,l}^{(p)}$ $:=\sum_{t\in E_{2(k-1)+l}^{(p-1)}}\mathbb{E}_{r\sim\hat{p}_{t}}\left[(\ell_{t}(\hat{y}_{t})-\ell_{t}(\mathcal{A}_{l,r}(t)))^{2}\right]$
$=\sum_{t\in E_{2(k-1)+l}^{(p-1)}}\frac{\sum_{r\in[r_{l}^{(p-1)}]}e^{\eta_{t}R_{r^{\prime},t-1}}(\ell_{t}(\hat{y}_{t})-\ell_{t}(\mathcal{A}_{l,r^{\prime}}(t)))^{2}}{\sum_{r^{\prime}\in[r_{l}^{(p-1)}]}e^{\eta_{t}R_{r^{\prime},t-1}}},$
where by abuse of notation we kept $R_{r^{\prime},t}$ to denote the cumulative regret compared to algorithm $r^{\prime}$ up to time $t$ during epoch $E_{2(k-1)+l}^{(p-1)}$. For convenience, let us denote by $Reg_{k,l}^{(p)}$ the regret incurred by the exponentially weighted algorithm A-Exp on each epoch $E_{2(k-1)+l}^{(p-1)}$ for $l\in[2]$ (see line 7 of Algorithm 4). Then,
$\sum_{t\in E_{k}^{(p)}}\ell_{t}(\textsc{R-Cover}_{k}^{(p)}(t))=\sum_{l\in[2]}\sum_{t\in E_{2(k-1)+l}^{(p-1)}}\ell_{t}(\textsc{R-Cover}_{k}^{(p)}(t))\\ \leq\sum_{l\in[2]}\min\left\{\sum_{t\in E_{2(k-1)+l}^{(p-1)}}\ell_{t}\left(f_{k,S}^{(p)}(x_{t})\right),\sum_{t\in E_{2(k-1)+l}^{(p-1)}}\ell_{t}\left(\textsc{R-Cover}_{2(k-1)+l}^{(p-1)}(t)\right)\right\}+\sum_{l\in[2]}Reg_{k,l}^{(p)}.$ (12)
[…]
$\sum_{l\in[2]}Reg_{k,l}^{(p)}$ $\leq\sum_{l\in[2]}12\sqrt{\max\left(\tilde{\Delta}_{k,l}^{(p-1)},1\right)\ln\left(\mathcal{N}(\mathcal{F};\epsilon,T)+1\right)}+4\ln\frac{2}{\delta}.$
Instead of working with the quantities $\tilde{\Delta}_{k,l}^{(p)}$, we instead define
$\Delta_{k}^{(p)}:=\Delta_{k,1}^{(p)}+\Delta_{k,2}^{(p)},\quad p\in[P],k\in[N_{p}].$
Applying Jensen’s inequality gives
$\sum_{l\in[2]}Reg_{k,l}^{(p)}\leq 12\sqrt{2\max\left(\Delta_{k}^{(p)},2\right)\ln\left(\mathcal{N}(\mathcal{F};\epsilon,T)+1\right)}+4\ln\frac{2}{\delta}.$ (13)
We are now ready to decompose the regret of the algorithm along the learning trajectory using the previous bound recursively. We start from the level $P$ and go down to some fixed depth $p_{0}\in\{0,\ldots,P\}$. Using Eq. 12 gives
$\sum_{t\in E_{1}^{(P)}}\ell_{t}(\hat{y}_{t})\leq\sum_{p=\max(p_{0},1)}^{P}\sum_{k\in[N_{p}],l\in[2]}Reg_{k,l}^{(p)}+\sum_{k\in[N_{p_{0}}]}\sum_{t\in E_{k}^{(p_{0})}}\ell_{t}\left(f_{k,S}^{(p_{0})}(x_{t})\right)$
Next, using Eq. 13, with probability at least $1-\delta\sum_{p=p_{0}}^{P}N_{p}\leq 1-2N_{p_{0}}\delta T\leq 1-2\delta T$ (recall that $T\geq 2^{P}$) we have for any choice of $p_{0}\in\{0,\ldots,P\}$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))\leq 12\sum_{p=\max(p_{0},1)}^{P}\sum_{k\in[N_{p}]}\sqrt{2\max\left(\Delta_{k}^{(p)},2\right)\ln\left(\mathcal{N}(\mathcal{F};\epsilon,T)+1\right)}\\ +8N_{p_{0}}\ln\frac{2}{\delta}+\sum_{k\in[N_{p_{0}}]}\sum_{t\in E_{k}^{(p_{0})}}\ell_{t}\left(f_{k,S}^{(p_{0})}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t})).$ (14)
Up to the last layer for $p=p_{0}$, the previous inequality shows that the regret of the algorithm essentially only corresponds to the regret accumulated by the learning with expert algorithms along the trajectory for the benchmark function $f^{\star}$. For convenience, we introduce for all $p\in\{0,\ldots,P\}$ and $k\in[N_{p}]$ the quantity
$\Lambda_{k}^{(p)}:=\sum_{t\in E_{k}^{(p)}}\ell_{t}\left(f_{k,S}^{(p)}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t})).$
We used a similar notation for $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$ for $p\in[P]$ because these terms will be bounded with the same techniques.
5.3 Bounding the regret term for each depth
We next bound each term of the right-hand side of Eq. 14 separately for each layer $p\in\{p_{0},\ldots,P\}$. That is, we need to bound the error terms $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$ for $k\in[N_{p}]$. Fix $p\in\{0,\ldots,P\}$ and $k\in[N_{p}]$ and let
$\mathcal{P}_{k}^{(p)}:=\left\{f\in\mathcal{F}:\max_{t\in[T_{k-1}^{(p)}]}\left|f(x_{t})-f^{\star}(x_{t})\right|\leq 2\epsilon\right\}.$ (15)
We then define for any $r\geq 1$,
$\Gamma_{k}^{(p,r)}$ $:=\sum_{t\in E_{k}^{(p)}}\gamma^{(p,r)}(t)\quad\text{where}\quad\gamma^{(p,r)}(t):=\sup_{f\in\mathcal{P}_{k}^{(p)}}\mathbb{E}\left[|f(x_{t})-f^{\star}(x_{t})|^{r}\mid\mathcal{H}_{t-1}\right],\quad t\in E_{k}^{(p)},$ (16)
Intuitively, $\Gamma_{k}^{(p,r)}$ quantifies the $\ell_{r}$ discrepancy between the queries on epoch $E_{k}^{(p)}$ and queries prior to this epoch. This measures the level of non-stationarity of the smooth process $(x_{t})_{t\in[T]}$ on each epoch. The following results shows that it suffices to bound $\Gamma_{k}^{(p,r)}$ to bound $\Delta_{k}^{(p)}$ and $\Lambda_{k}^{(p)}$.
Lemma 12.
With probability at least $1-\delta$,
$\Delta_{k}^{(p)}\leq 5\Gamma_{k}^{(p,2)}+16\ln\frac{T}{\delta},\quad p\in[P],k\in[N_{p}].$
Similarly, for any $p\in\{0,\ldots,P\}$, with probability at least $1-\delta$,
$\Lambda_{k}^{(p)}\leq 2\Gamma_{k}^{(p,1)}+3\ln\frac{T}{\delta},\quad k\in[N_{p}].$
Proof  Fix $p\in[P]$ and $k\in[N_{p}]$. During its run on epoch $E_{k}^{(p)}$, the learning with expert prediction algorithm A-Exp uses predictions from depth-$(p-1)$ algorithms. In practice, all considered sub-algorithms—that is, for epochs $E_{k^{\prime}}^{p^{\prime}}$ with $p^{\prime}<p$ and such that $E_{k^{\prime}}^{(p^{\prime})}\subset E_{k}^{(p)}$—are proper in the sense that they proceed by first selecting some predictor function $\hat{f}_{t}\in\mathcal{F}$ then implementing its prediction $\hat{f}_{t}(x_{t})$. The choice of the function $\hat{f}_{t}$ is randomized, but is made before observing the value $x_{t}$. As an important remark, all these potentially-selected functions belong to $\mathcal{F}(S_{k}^{(p)})$ since for sub-algorithms we append reference functions $(f_{i},t_{i})$ to the reference set $S_{k}^{(p)}$. Next, note that by construction, we have $(f_{k}^{(p)},T_{k-1}^{(p)})\in S_{k}^{(p)}$ (see the recursion line 7 of Algorithm 4). In particular, all these functions belong to $\mathcal{F}(S_{k}^{(p)})\subset B_{f_{k}^{(p)}}(\mathcal{F};\epsilon,T_{k-1}^{(p)}):=B_{k}^{(p)}$, where we introduced the last notation for simplicity. For $p=P$, we simply have $B_{1}^{(P)}=\mathcal{F}$.
[…]
$\Delta_{k}^{(p)}=\sum_{l\in[2]}\tilde{\Delta}_{k,l}^{(p)}$
$=\sum_{l\in[2]}\sum_{t\in E_{C(k-1)+l}^{(p-1)}}\mathbb{E}_{\hat{r}_{t}^{\prime}}\left[\left(\ell_{t}(f_{l,\hat{r}_{t},t}(x_{t}))-\ell_{t}(f_{l,\hat{r}_{t}^{\prime},t}(x_{t}))\right)^{2}\mid\mathcal{H}_{t},\,\hat{r}_{t},\,f_{r,l^{\prime},t},l^{\prime}\in[r_{l}^{(p-1)}]\right]$
$\leq\sum_{l\in[2]}\sum_{t\in E_{C(k-1)+l}^{(p-1)}}\mathbb{E}_{\hat{r}_{t}^{\prime}}\left[\left(f_{l,\hat{r}_{t},t}(x_{t})-f_{l,\hat{r}_{t}^{\prime},t}(x_{t})\right)^{2}\mid\mathcal{H}_{t},\,\hat{r}_{t},\,f_{r,l^{\prime},t},l^{\prime}\in[r_{l}^{(p-1)}]\right]$
$\leq 2\sum_{l\in[2]}\sum_{t\in E_{C(k-1)+l}^{(p-1)}}\underbrace{\mathbb{E}_{\hat{r}_{t}^{\prime}}\left[\left(f_{l,\hat{r}_{t},t}(x_{t})-f^{\star}(x_{t})\right)^{2}+\left(f_{l,\hat{r}_{t}^{\prime},t}(x_{t})-f^{\star}(x_{t})\right)^{2}\mid\mathcal{H}_{t},\,\hat{r}_{t},\,f_{r,l^{\prime},t},l^{\prime}\in[r_{l}^{(p-1)}]\right]}_{X_{t}^{(p)}},$
where we used the identity $(a+b)^{2}\leq 2(a^{2}+b^{2})$ for $a,b\geq 1$. Next, note that
$Y_{t}^{(p)}$ $:=\mathbb{E}_{x_{t},\hat{r}_{t}}\left[X_{t}^{(p)}\mid\mathcal{H}_{t-1},\,f_{r,l^{\prime},t},l^{\prime}\in[r_{l}^{(p-1)}]\right]$
$=\mathbb{E}_{\hat{r}_{t},\hat{r}_{t}^{\prime}}\left[\mathbb{E}_{x_{t}\mid\mathcal{H}_{t-1}}\left[\left(f_{l,\hat{r}_{t},t}(x_{t})-f^{\star}(x_{t})\right)^{2}+\left(f_{l,\hat{r}_{t}^{\prime},t}(x_{t})-f^{\star}(x_{t})\right)^{2}\mid\mathcal{H}_{t-1},\,f_{l,\hat{r}_{t},t},\,f_{l,\hat{r}_{t}^{\prime},t}\right]\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.\mid\mathcal{H}_{t-1},\,f_{r,l^{\prime},t},l^{\prime}\in[r_{l}^{(p-1)}]\right]$
$\leq 2\sup_{f\in B_{k}^{(p)}}\mathbb{E}_{x_{t}\mid\mathcal{H}_{t-1}}\left[\left(f(x_{t})-f^{\star}(x_{t})\right)^{2}\mid\mathcal{H}_{t-1}\right]$
$\leq 2\gamma^{(p,2)}(t).$
In the last inequality, we used the definition of $B_{k}^{(p)}=B_{f_{k}^{(p)}}(\mathcal{F};\epsilon,T_{k-1}^{(p)})$ together with the fact that by construction $f^{\star}\in\mathcal{F}(S_{k}^{(p)})\subset B_{k}^{(p)}$. Hence, the triangle inequality implies that $B_{k}^{(p)}\subset\mathcal{P}_{k}^{(p)}$.
The previous equation shows that $\Delta_{k}^{(p)}-4\Gamma_{k}^{(p,2)}\leq 2\sum_{t\in E_{p}^{(k)}}(X_{t}^{(p)}-2\gamma^{(p,2)}(t))$ where the right-hand side is a sum of super-martingale differences. Further, these differences are bounded in absolute value by $|X_{t}^{(p)}-2\gamma^{(p,2)}(t)|\leq 4$. To bound $\Delta_{k}^{(p)}$ in terms of $\Gamma_{k}^{(p,2)}$, we can directly use Azuma-Hoeffding’s inequality which would give an extra term of the form $\sqrt{(T_{k}^{(p)}-T_{k-1}^{(p)})\ln\frac{1}{\delta}}$ for a bound with probability $1-\delta$. Because we will consider cases for which $\Gamma_{k}^{(p)}$ is significantly smaller than $\sqrt{T_{k}^{(p)}-T_{k-1}^{(p)}}$, we instead use Freedman’s inequality stated in Lemma 19 to the sum $\sum_{t\in E_{p}^{(k)}}X_{t}^{(p)}-Y_{t}^{(p)}$ using the filtration $\mathcal{F}_{t}=\sigma(X_{t^{\prime}}^{(p)},t^{\prime}\leq t;Y_{t^{\prime}}^{(p)},t^{\prime}\leq t+1)$ (note that $\mathbb{E}[X_{t}^{(p)}\mid\mathcal{F}_{t-1}]=Y_{t}^{(p)}$). To do so, we compute
$\sum_{t\in E_{k}^{(p)}}\mathbb{E}\left[(X_{t}^{(p)}-Y_{t}^{(p)})^{2}\mid\mathcal{F}_{t-1}\right]$ $\leq\sum_{t\in E_{k}^{(p)}}\mathbb{E}\left[(X_{t}^{(p)})^{2}\mid\mathcal{F}_{t-1}\right]$
$\overset{(i)}{\leq}2\sum_{t\in E_{k}^{(p)}}\mathbb{E}\left[X_{t}^{(p)}\mid\mathcal{F}_{t-1}\right]=2\sum_{t\in E_{k}^{(p)}}Y_{t}^{(p)}$
$\leq 4\sum_{t\in E_{k}^{(p)}}\gamma^{(p,2)}(t)=4\Gamma_{k}^{(p,2)}.$
In $(i)$ we used the fact that $|X_{t}^{(p)}|\leq 2$ since functions in $\mathcal{F}$ have value in $[0,1]$. Last, we always have $|X_{t}^{(p)}-Y_{t}^{(p)}|\leq 4$. Then, Lemma 19 with the union bound over all $p\in[P]$ and $k\in[N_{p}]$ implies that with probability at least $1-\delta$, using $\eta=1/8$,
$\Delta_{k}^{(p)}\leq 2\sum_{t\in E_{k}^{(p)}}(X_{t}^{(p)}-Y_{t}^{(p)})+4\Gamma_{k}^{(p,2)}\leq 5\Gamma_{k}^{(p,2)}+16\ln\frac{T}{\delta},\quad p\in[P],k\in[N_{p}].$ (17)
Here we used the fact that $\sum_{p=1}^{P}N_{p}=\sum_{p=1}^{P}2^{P-p}\leq T$.
We next bound the terms $\Lambda_{k}^{(p)}$ in a similar fashion for a fixed $p\in\{0,\ldots,P\}$. First, note that by construction, for any $k\in[N_{p}]$, we still have
$f_{k,S}^{(p)},f^{\star}\in\mathcal{F}(S_{k}^{(p)})\subset B_{k}^{(p)}:=B_{f_{k}^{(p)}}(\mathcal{F};\epsilon,T_{k-1}^{(p)}).$
As a result, as discussed above $f_{k,S}^{(p)}\in\mathcal{P}_{k}^{(p)}$. Further, $f_{k,S}^{(p)}$ is fixed at the beginning of epoch $E_{k}^{(p)}$ hence can be made $\mathcal{H}_{T_{k-1}^{(p)}}$-measurable without loss of generality. Then, for any $k\in[N_{p}]$ using the fact that the losses are $1$-Lipschitz,
$\Lambda_{k}^{(p)}\leq\sum_{t\in E_{k}^{(p)}}\underbrace{\left|f_{k,S}^{(p)}(x_{t})-f^{\star}(x_{t})\right|}_{\tilde{X}_{t}^{(p)}}$
and similarly as before,
$\tilde{Y}_{t}^{(p)}$ $:=\mathbb{E}_{x_{t},\hat{r}_{t}}\left[X_{t}^{(p)}\mid\mathcal{H}_{t-1},\,f_{k,S}^{(p)}\right]$
$=\mathbb{E}_{\hat{r}_{t},\hat{r}_{t}^{\prime}}\left[\mathbb{E}_{x_{t}\mid\mathcal{H}_{t-1}}\left[\left|f_{k,S}^{(p)}(x_{t})-f^{\star}(x_{t})\right|\mid\mathcal{H}_{t-1},\,f_{k,S}^{(p)}\right]\mid\mathcal{H}_{t-1},\,f_{k,S}^{(p)}\right]\leq\gamma^{(p,1)}(t).$
We can also bound $|\tilde{X}_{t}^{(p)}-\tilde{Y}_{t}^{(p)}|\leq 2$. Again, we use Freedman’s inequality to $\sum_{t\in E_{k}^{(p)}}\tilde{X}_{t}^{(p)}-\tilde{Y}_{t}^{(p)}$ noting that
$\sum_{t\in E_{p}^{(k)}}\mathbb{E}\left[(\tilde{X}_{t}^{(p)}-\tilde{Y}_{t}^{(p)})^{2}\mid\mathcal{F}_{t-1}\right]\leq\sum_{t\in E_{p}^{(k)}}\mathbb{E}\left[(\tilde{X}_{t}^{(p)})^{2}\mid\mathcal{F}_{t-1}\right]\leq 2\sum_{t\in E_{p}^{(k)}}\mathbb{E}\left[\tilde{X}_{t}^{(p)}\mid\mathcal{F}_{t-1}\right]\leq 2\Gamma_{k}^{(p,1)}.$
Similarly as before, Lemma 19 with $\eta=1/3$ with the union bound on all $k\in[N_{p}]$ then implies that with probability at least $1-\delta$,
$\Lambda_{k}^{(p)}\leq\sum_{t=1}^{T}(\tilde{X}_{t}^{(p)}-\tilde{Y}_{t}^{(p)})+\Gamma_{k}^{(p,1)}\leq 2\Gamma_{k}^{(p,1)}+3\ln\frac{T}{\delta},\quad k\in[N_{p}].$
Here we used $N_{p}\leq T$. This ends the proof of the lemma. $\blacksquare$
We denote by $L_{p}:=\left\lfloor T/N_{p}\right\rfloor+1$ the maximum length of each depth-$p$ epoch. We recall that from Eq. 10 the depth-$p$ epochs all have length $L_{p}$ of $L_{p}-1$. Note that by construction because $p\geq p_{0}\geq 1$, we have $L_{p}\geq 2$ and hence $L_{p}-1\geq 1$. In the worst case, each term $\Gamma_{k}^{(p)}$ for $k\in[N_{p}]$ could be as large as $L_{p}$. We show, however, that smoothness ensures that such epochs are very few.
Lemma 13.
Fix $r\geq 1$, $p\in\{0,\ldots,P\}$ and suppose that $(x_{t})_{t\in[T]}$ is a $\sigma$-smooth stochastic process with respect to some measure $\mu$, where $T\geq 2$. For any parameters $w\geq 2$ and $q\in(0,1]$ satisfying
$q\geq 12\sqrt{(2\epsilon)^{r}\frac{2\ln(eT)}{\sigma}},$ (18)
with probability at least $1-\delta$,
|$\left|\left\{k\in[N_{p}]:\sum_{t\in E_{k}^{(p)}}\gamma^{(p,r)}(t)\mathbbm{1}[\gamma^{(p,r)}(t)\geq q]\geq w\right\}\right|\\ \leq\frac{c_{0}r\ln^{2}T}{q\sigma w}\left(\ln\mathbb{E}_{\mu}\left[W_{8\frac{T}{\sigma}\ln(\frac{T}{2\delta})}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w\right),$
for some universal constant $c_{0}>0$. In particular, if
$q\geq\max\left(24\sqrt{(2\epsilon)^{r}\frac{2\ln(eT)}{\sigma}},\frac{2w}{L_{p}-1}\right),$ (19)
then with probability at least $1-\delta$,
|$\left|\left\{k\in[N_{p}]:\Gamma_{k}^{(p,r)}\geq q(T_{k}^{(p)}-T_{k-1}^{(p)})\right\}\right|\leq\frac{2c_{0}r\ln^{2}T}{q\sigma w}\left(\ln\mathbb{E}_{\mu}\left[W_{8\frac{T}{\sigma}\ln(\frac{T}{2\delta})}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w\right).$
We defer the proof of this result to Section 5.5. We now select the parameter
$w=w(T,\delta):=\ln\mathbb{E}_{\mu}\left[W_{8T\ln(\frac{T}{2\delta})/\sigma}\left(\mathcal{F}\right)\right]+10\ln\frac{T}{\delta}+2$
which satisfies $w\geq 2$. We combine Lemmas 12 and 13 both for the probability tolerance $\delta$, which implies that for any $w\geq 2$ and $q\in(0,1]$ satisfying Eq. 19 for $r=2$, with probability at least $1-2\delta$,
|$\left|\left\{k\in[N_{p}]:\Delta_{k}^{(p)}\geq 6q(T_{k}^{(p)}-T_{k-1}^{(p)})\right\}\right|$
$\overset{(i)}{\leq}\left|\left\{k\in[N_{p}]:\Delta_{k}^{(p)}\geq 5q(T_{k}^{(p)}-T_{k-1}^{(p)})+20\ln\frac{T}{\delta}\right\}\right|$
$\overset{(ii)}{\leq}\left|\left\{k\in[N_{p}]:\Gamma_{k}^{(p,2)}\geq q(T_{k}^{(p)}-T_{k-1}^{(p)})\right\}\right|$
$\overset{(iii)}{\leq}\frac{c_{0}\ln^{2}T}{q\sigma w(T,\delta)}\left(\ln\mathbb{E}_{\mu}\left[W_{8T\ln(\frac{T}{2\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w(T,\delta)\right)\leq\frac{2c_{0}\ln^{2}T}{q\sigma},$ (20)
for some constant $c_{0}\geq 1$. In $(i)$ we used the fact that $q\geq\frac{2w}{L_{p}-1}\geq\frac{20\ln\frac{T}{\delta}}{T_{k}^{(p)}-T_{k-1}^{(p)}}$. In $(ii)$ we used Lemma 12 and in $(iii)$ we used Lemma 13. Similarly, for any $w\geq 2$ and $q\in(0,1]$ satisfying Eq. 19 for $r=1$, with probability at least $1-2\delta$,
|$\left|\left\{k\in[N_{p}]:\Lambda_{k}^{(p)}\geq 3q(T_{k}^{(p)}-T_{k-1}^{(p)})\right\}\right|\leq\frac{2c_{0}\ln^{2}(2T)}{q\sigma}.$ (21)
Bounding the regret term involving $\Delta_{k}^{(p)}$.
Using Eqs. 20 and 21, we can now bound the regret terms from the decomposition in Eq. 14. We start with the term involving the quantities $\Delta_{k}^{(p)}$. For $p\in\{p_{0},\ldots,P\}$ we let
$q_{0}^{(p)}:=300\cdot\max\left(2\epsilon\sqrt{\frac{\ln T}{\sigma}},\frac{w(T,\delta)}{L_{p}-1},\frac{c_{0}\ln^{2}T}{\sigma N_{p}}\right).$
If $q_{0}^{(p)}\geq 1$, we can simply bound
$\sum_{k\in[N_{p}]}\sqrt{\max\left(\Delta_{k}^{(p)},2\right)}\overset{(i)}{\leq}\sum_{k\in[N_{p}]}2\sqrt{T_{k}^{(p)}-T_{k-1}^{(p)}}\overset{(ii)}{\leq}2\sqrt{TN_{p}}\leq 2\sqrt{q_{0}^{(p)}TN_{p}}.$ (22)
In $(i)$ we used the fact that $\Delta_{k}^{(p)}$ is a sum of terms bounded by $1$ since the loss is $1$-Lipschitz and the functions $f\in\mathcal{F}$ have value within $[0,1]$. In $(ii)$ we used Jensen’s inequality.
Otherwise, if $q_{0}^{(p)}\leq 1$, we introduce the parameters $q_{s}^{(p)}=4^{s}q_{0}^{(p)}$ for $s\geq 0$ and let $s_{p}$ be the last index for which $q_{s}^{(p)}\leq 4$. We then define the sets
$\mathcal{T}^{(p)}(s):=\left\{k\in[N_{p}]:q_{s}^{(p)}(T_{k}^{(p)}-T_{k-1}^{(p)})\leq\Delta_{k}^{(p)}<q_{s+1}^{(p)}(T_{k}^{(p)}-T_{k-1}^{(p)})\right\},\quad s\in\{0,\ldots,s_{p}\}.$
By construction of $s_{p}$, any epoch $k\in[N_{p}]$ either belongs to one of the sets above or satisfies $\Delta_{k}^{(p)}\leq q_{0}^{(p)}(T_{k}^{(p)}-T_{k-1}^{(p)})$. Also, note that there exists a constant $c>0$ such that $s_{p}\leq c\ln T$ since $L_{p},N_{p}\leq T$. We also recall that $P\leq\log_{2}(T)$. Hence, up to changing the constant $c>0$, Eq. 20 implies that on some event $\mathcal{E}_{\delta}$ with probability at least $1-c\delta\ln^{2}T$, for all $p\in\{p_{0},\ldots,P\}$ such that $q_{0}^{(p)}\leq 1$,
$\left|\mathcal{T}^{(p)}(s)\right|\leq\frac{12c_{0}\ln^{2}T}{q_{s}^{(p)}\sigma},\quad s\in\{0,\ldots,s_{p}\}.$
Hence, on $\mathcal{E}$, for any $p\in\{p_{0},\ldots,P\}$ such that $q_{0}^{(p)}\leq 1$, we have
$\sum_{k\in[N_{p}]}\sqrt{\max\left(\Delta_{k}^{(p)},2\right)}$ $\overset{(i)}{\leq}\sum_{k\in[N_{p}]}\sqrt{q_{0}^{(p)}(T_{k}^{(p)}-T_{k-1}^{(p)})}+\sum_{s=0}^{s_{p}}|\mathcal{T}^{(p)}(s)|\cdot\sqrt{q_{s+1}^{(p)}L_{p}}$
$\overset{(ii)}{\leq}\sqrt{q_{0}^{(p)}TN_{p}}+\frac{12c_{0}\ln^{2}T}{\sigma}\sum_{s=0}^{s_{p}}\frac{\sqrt{q_{s+1}^{(p)}L_{p}}}{q_{s}^{(p)}}$
$=\sqrt{q_{0}^{(p)}TN_{p}}+\frac{24c_{0}\ln^{2}T}{\sigma}\sum_{s=0}^{s_{p}}\sqrt{\frac{L_{p}}{q_{s}^{(p)}}}$
$\overset{(iii)}{\leq}\sqrt{q_{0}^{(p)}TN_{p}}+\frac{48c_{0}\ln^{2}T}{\sigma}\sqrt{\frac{2T}{q_{0}^{(p)}N_{p}}}$
$\overset{(iv)}{\leq}2\sqrt{q_{0}^{(p)}TN_{p}}.$ (23)
In $(i)$ we used the fact that $q_{0}^{(p)}(T_{k}^{(p)}-T_{k-1}^{(p)})\geq 2w\geq 2$ to delete the maximum with $2$, and in $(ii)$ we used Jensen’s inequality. In $(iii)$ we use the fact that $L_{p}N_{p}\leq T+N_{p}\leq 2T$ and in $(iv)$ we used the fact that $q_{0}^{(p)}\geq 48\sqrt{2}\frac{c_{0}\ln^{2}T}{\sigma N_{p}}$.
As a result, on the event $\mathcal{E}_{\delta}$, we can combine Eqs. 22 and 23 to obtain
$\sum_{p=p_{0}}^{P}\sum_{k\in[N_{p}]}\sqrt{\max\left(\Delta_{k}^{(p)},2\right)}$ $\lesssim\sqrt{\epsilon}\left(\frac{\ln T}{\sigma}\right)^{1/4}\sum_{p=p_{0}}^{P}\sqrt{TN_{p}}+\sqrt{w(T,\delta)}\sum_{p=p_{0}}^{P}\sqrt{\frac{TN_{p}}{L_{p}-1}}$
$\qquad\qquad+(P-p_{0}+1)\sqrt{\frac{\ln^{2}T}{\sigma}\cdot T}$
$\lesssim\sqrt{\epsilon TN_{p_{0}}}+N_{p_{0}}\sqrt{w(T,\delta)}+\ln^{2}T\sqrt{\frac{T}{\sigma}},$ (24)
where the $\lesssim$ symbol only hides universal constants.
Bounding the regret term involving $\Lambda_{k}^{(p)}$.
We next turn to last regret term from the decomposition in Eq. 14. We let
$\tilde{q}_{0}^{(p)}:=300\cdot\max\left(\sqrt{2\epsilon\frac{\ln T}{\sigma}},\frac{w(T,\delta)}{L_{p}-1},\frac{c_{0}\ln^{3}T}{\sigma N_{p}}\right).$
If $\tilde{q}_{0}^{(p_{0})}\geq 2$, the resulting regret bound is vacuous. Hence, we focus on the case when $\tilde{q}_{0}^{(p_{0})}\geq 2$. As before, we let $\tilde{q}_{s}^{(p_{0})}=2^{s}\tilde{q}_{0}^{(p_{0})}$ for $s\geq 0$ and let $\tilde{s}_{p_{0}}$ be the last index such that $\tilde{q}_{s}^{(p_{0})}\leq 2$. As above, we have $\tilde{s}_{p_{0}}\leq c\ln T$ for some constant $c>0$. Then, Eq. 21 implies that on an event $\mathcal{F}_{\delta}$ of probability at least $1-c\delta\ln T$, for all $s\in\{0,\ldots,\tilde{s}_{p_{0}}\}$, we have
|$\left|\left\{k\in[N_{p}]:\tilde{q}_{s}^{(p_{0})}(T_{k}^{(p_{0})}-T_{k-1}^{(p_{0})})\leq\Lambda_{k}^{(p_{0})}<\tilde{q}_{s+1}^{(p_{0})}(T_{k}^{(p_{0})}-T_{k-1}^{(p_{0})})\right\}\right|\leq\frac{6c_{0}\ln^{2}T}{\tilde{q}_{s}^{(p_{0})}\sigma}.$
Using the same arguments as above shows that on $\mathcal{F}_{\delta}$,
$\sum_{k\in[N_{p_{0}}]}\Lambda_{k}^{(p_{0})}$ $\leq\tilde{q}_{0}^{(p_{0})}T+\frac{6c_{0}\ln^{2}T}{\sigma}\sum_{s=0}^{\tilde{s}_{p_{0}}}\frac{\tilde{q}_{s+1}^{(p_{0})}L_{p_{0}}}{\tilde{q}_{s}^{(p_{0})}}$
$\leq\tilde{q}_{0}^{(p_{0})}T+\frac{12c_{0}\ln^{2}T}{\sigma}(\tilde{s}_{p_{0}}+1)L_{p_{0}}$
$\leq\tilde{q}_{0}^{(p_{0})}T+\frac{24cc_{0}\ln^{3}T}{\sigma}L_{p_{0}}$
$\leq(1+c)\tilde{q}_{0}^{(p_{0})}T.$ (25)
In the last inequality, we used $N_{p_{0}}L_{p_{0}}\leq 2T$ and the definition of $\tilde{q}_{0}^{(p_{0})}$. Note that Eq. 25 also trivially holds if $\tilde{q}_{0}^{(p_{0})}\geq 2$.
Final regret bound
We now combine the bounds from Eqs. 24 and 25 within the regret decomposition from Eq. 14 which shows that on $\mathcal{E}_{\delta}\cap\mathcal{F}_{\delta}$ of probability at least $1-2c\delta\ln^{2}T$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))$ $\lesssim\left(\sqrt{\epsilon TN_{p_{0}}}+N_{p_{0}}\sqrt{w(T,\delta)}+\ln^{2}T\sqrt{\frac{T}{\sigma}}\right)\sqrt{\ln\left(\mathcal{N}(\mathcal{F};\epsilon,T)+1\right)}$
$\qquad\qquad+\sqrt{\frac{\epsilon\ln T}{\sigma}}\cdot T+w(T,\delta)N_{p_{0}}+\frac{\ln^{3}T}{\sigma N_{p_{0}}}T.$ (26)
Here we used the fact that $w(T,\delta)\geq\ln\frac{1}{\delta}$ to delete the term $N_{p_{0}}\ln\frac{1}{\delta}$. This holds for all $p_{0}\in[P]$. Hence, we obtain the following result which implies in particular Theorem 6.
Theorem 14.
Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class with VC dimension $d$. Suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. Then, R-Cover (Recursive Covering) with the parameter $\epsilon\in[0,1]$ makes predictions $\hat{y}_{t}$ such that for any function $f^{\star}\in\mathcal{F}$, with probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-$ $\sum_{t=1}^{T}\ell_{t}(f^{\star}(x_{t}))$
$\lesssim\min_{N_{0}}\left\{\left(\sqrt{\epsilon TN_{0}}+N_{0}\sqrt{w(T,\delta)}+\ln^{2}T\sqrt{\frac{T}{\sigma}}\right)\sqrt{\ln\left(\mathcal{N}(\mathcal{F};\epsilon,T)+1\right)}\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\sqrt{\frac{\epsilon\ln T}{\sigma}}\cdot T+N_{0}w(T,\delta)+\frac{\ln^{3}T}{\sigma N_{0}}T\right\},$ (27)
where
$w(T,\delta)=\ln\mathbb{E}_{\mu}\left[W_{8T\ln(\frac{cT\ln^{2}T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta},$
for some universal constant $c>0$. we recall that the covering numbers can be bounded in terms of the fat-shattering dimension via Theorem 2.
For instance, if there exists $d\geq 1$ such that $\mathsf{fat}_{\mathcal{F}}(r)\leq d\ln\frac{1}{r}$ for all $r>0$, the regret bound for R-Cover with parameter $\epsilon=1/T$ becomes
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\sum_{t=1}^{T}\ell_{t}(f^{\star}(x_{t}))\leq C\sqrt{\frac{\left(d\ln^{3}\left(T\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}\right)\ln^{3}T}{\sigma}\cdot T},$
for some constant $C>0$.
If $\mathsf{fat}_{\mathcal{F}}(r)\lesssim r^{-p}$ for $p>0$, the regret bound for R-Cover with parameter $\epsilon=\left(\frac{\ln T}{T}\right)^{\frac{1}{p+1}}$ becomes
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\sum_{t=1}^{T}\ell_{t}(f^{\star}(x_{t}))\lesssim_{p}\frac{\ln^{3}T}{\sqrt{\sigma}}\cdot T^{1-\frac{1}{2(p+1)}}+\frac{\ln^{4}T\ln^{3}\frac{T}{\delta}}{\sigma}\cdot T^{1-\frac{1}{2(p+1)}-\frac{\min(p,1)}{2(p+1)(p+2)}}.$
where $\lesssim_{p}$ only hides factors depending (potentially exponentially) in $p$.
Proof  Eq. 26 that Eq. 27 directly holds if the minimum is taken over $N_{0}\in\{N_{p_{0}}=2^{P-p_{0}},p_{0}\in[P]\}$. Hence, up to a factor of two, the regret bound holds if the minimum is taken over $N_{0}\in[T]$. We next observe that for $N_{0}\gtrsim T$ or $N_{0}\lesssim 1$, the bound exceeds $2T$, hence trivially holds.
We now turn to the next two claims. Observe that in both cases, if $\sigma\lesssim\frac{1}{T}$, the bound trivially holds. Without loss of generality, we therefore suppose that $\sigma\gtrsim\frac{1}{T}$ from now.
When $\mathsf{fat}_{\mathcal{F}}(r)\leq d\ln\frac{1}{r}$ for $r>0$, Theorem 2 with $\alpha\lesssim\frac{1}{\ln\ln T}$ then implies that for all $\epsilon\in[\frac{1}{T^{2}},\frac{1}{\sqrt{T}}]$,
$\ln(\mathcal{N}(\mathcal{F};\epsilon,T)+1)\lesssim\mathsf{fat}_{\mathcal{F}}(c\alpha\epsilon)\ln^{1+\alpha}\frac{T}{\epsilon}\lesssim d\ln^{2}T.$
We recall that we assumed $\sigma\gtrsim 1/T$. Similarly, given the target bound, if $d\gtrsim T$ the result is immediate. We also suppose that $d\lesssim T$ from now. Next, by Proposition 18, we have
$w(T,\delta)\lesssim d\ln^{3}\left(\frac{T}{\sigma}\ln\frac{T}{\delta}\right)+\ln\frac{T}{\delta}\lesssim d\ln^{3}\left(T\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}.$
We then choose the parameter $\epsilon=1/T$ and the value $N_{0}=\sqrt{\frac{\ln^{3}T}{\sigma\left(d\ln^{3}\left(T\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}\right)}\cdot T}$, which gives the desired bound.
We next turn to the case when $\mathsf{fat}_{\mathcal{F}}(r)\lesssim r^{-p}$. In the rest of the proof, the symbols $\lesssim_{p}$ may hide factors in $p$ of the form $c^{p}$ for universal constants. In this case, Theorem 2 implies that for $\epsilon\in[\frac{1}{T^{2}},1]$,
$\ln(\mathcal{N}(\mathcal{F};\epsilon,T)+1)\lesssim\frac{\ln^{2}T}{\epsilon^{p}}.$
Hence, by Proposition 18,
$w(T;\delta)\lesssim_{p}\left(\frac{T}{\sigma}\right)^{\alpha(p)}\ln^{3}\frac{T}{\delta},\quad\text{where}\quad\alpha(p):=\begin{cases}\frac{p}{2+p}&0<p\leq 2\\ 1-\frac{1}{p}&p\geq 2.\end{cases}$
Again, here we used $\sigma\gtrsim 1/T$. For intuition, the two main terms in the regret bound for Eq. 27 are $\ln^{2}T\sqrt{T\ln(\mathcal{N}(\mathcal{F};\epsilon,T)+1)/\sigma}$ and $T\sqrt{\epsilon\ln T/\sigma}$. To minimize these, we then choose the parameter $\epsilon=\left(\frac{\ln T}{T}\right)^{\frac{1}{p+1}}$. With $N_{0}=\frac{\ln^{3}T}{\sqrt{\sigma}}\cdot T^{\frac{1}{2(p+1)}}$ we obtain the following regret bound,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\sum_{t=1}^{T}\ell_{t}(f^{\star}(x_{t}))$ $\lesssim_{p}\frac{\ln^{3}T}{\sqrt{\sigma}}T^{1-\frac{1}{2(p+1)}}+\frac{\ln^{3}T\ln^{3}\frac{T}{\delta}}{\sigma^{\frac{1}{2}+\alpha(p)}}T^{\frac{1}{2(p+1)}+\alpha(p)}+\frac{\ln^{4}T\ln^{\frac{3}{2}}\frac{T}{\delta}}{\sigma^{\frac{\alpha(p)+1}{2}}}T^{\frac{1}{2}+\frac{\alpha(p)}{2}}.$
Together with $\alpha(p)\leq 1-\frac{1}{p+1}-\frac{\min(p,1)}{(p+1)(p+2)}$, this implies the desired bound. $\blacksquare$
As a remark, for function classes with finite VC dimension $d$, we can use the tighter bound on the Wills functional from Proposition 18 which gives $\ln W_{m}(\mathcal{F})\lesssim d\ln m$. Further, for VC classes, we can simply use $\epsilon=0$ since Sauer-Shela’s lemma (Lemma 1) guarantees $\ln\mathcal{N}(\mathcal{F};0,T)\lesssim d\ln T$. Altogether, this gives the following slightly improved bound.
Proposition 15.
Let $\mathcal{F}:\mathcal{X}\to\{0,1\}$ be a function class with VC dimension $d$. Suppose that $(x_{t})_{t\geq 1}$ is a $\sigma$-smooth sequence on $\mathcal{X}$ with respect to some unknown base measure $\mu$. Then, R-Cover with $\epsilon=0$ makes predictions $\hat{y}_{t}$ such that for any $f^{\star}\in\mathcal{F}$, with probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\sum_{t=1}^{T}\ell_{t}(f^{\star}(x_{t}))\leq C\sqrt{\frac{(d\ln^{2}T+d\ln\ln\frac{1}{\delta}+\ln\frac{1}{\delta})\ln^{3}T}{\sigma}\cdot T}.$
for some universal constant $C>0$.
5.4 From oblivious to adaptive benchmarks for classification
Theorem 14 gives high-probability bounds for the oblivious regret of R-Cover. In the specific case of classification, we can further extend these bounds to the adaptive regret of the algorithm. In this section, we therefore focus on the case where $\mathcal{F}:\mathcal{X}\to\{0,1\}$ is a function class with finite VC dimension $d$, using ideas inspired from [36].
First construct an $\epsilon$-cover $\mathcal{H}$ of the function class $\mathcal{F}$ for the base measure $\mu$. Formally, an $\epsilon$-cover is a subset of $\mathcal{F}$ such that for all $f\in\mathcal{F}$ there exists $h\in\mathcal{H}$ with
$\mathbb{P}_{x\sim\mu}(f(x)\neq h(x))\leq\epsilon.$
Since $\mathcal{F}$ has VC dimension $d$, we can ensure $\ln|\mathcal{H}|\leq 2d\ln(e^{2}/\epsilon)$ (see [32] or [12, Lemma 13.6]). Taking the union bound for all (non-adaptive) benchmark functions in $\mathcal{H}$, Proposition 15 implies that with probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{H}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))$ $\leq C\sqrt{\frac{(d\ln^{2}T+d\ln\ln\frac{|\mathcal{H}|}{\delta}+\ln\frac{|\mathcal{H}|}{\delta})\ln^{3}T}{\sigma}\cdot T}$
$\leq C^{\prime}\sqrt{\frac{(d\ln^{2}T+d\ln\ln\frac{1}{\epsilon\delta}+d\ln\frac{1}{\epsilon}+\ln\frac{1}{\delta})\ln^{3}T}{\sigma}\cdot T},$ (28)
for some constant $C^{\prime}\geq 1.$ In the last inequality, we used the fact that without loss of generality, $d\lesssim T$, otherwise the regret bound from Theorem 4 is immediate. Next, for any function $f\in\mathcal{F}$, denote by $h_{f}\in\mathcal{H}$ a function such that $\mathbb{P}_{\mu}(f\neq h)\leq\epsilon$. Then,
$\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\geq\sum_{t=1}^{T}\ell_{t}(h_{f}(x_{t}))-\sum_{t=1}^{T}\mathbbm{1}(f(x_{t})\neq h_{f}(x_{t})).$
As a result, denoting by $\mathcal{G}:=\{\mathbbm{1}[f\neq h_{f}],f\in\mathcal{F}\}$, we can decompose the adaptive regret via
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\leq\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{H}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))+\sup_{g\in\mathcal{G}}\sum_{t=1}^{T}g(x_{t}).$ (29)
[36, Lemma 3.3] directly bounds the expected value of $\sup_{g\in\mathcal{G}}\sum_{t=1}^{T}g(x_{t})$. Combined with Section 5.4, this already gives a bound for the expected adaptive regret. To give useful intuitions and get high-probability bounds, we detail the steps of the proof below.
Importantly, by construction of the $\epsilon$-cover, we have $\mathbb{E}_{x\sim\mu}[g(x)]\leq\epsilon$ for all $g\in\mathcal{G}$. Also, $\mathcal{G}$ has VC dimension at most $2d$. [36] then use a coupling argument to reduce to the i.i.d. case for which VC theory yields uniform convergence bounds using Lemma 11. On the event $\mathcal{E}_{k}$ from Lemma 11, we have
$\sup_{g\in\mathcal{G}}\sum_{t=1}^{T}g(x_{t})\leq\sup_{g\in\mathcal{G}}\sum_{t=1}^{T}\sum_{j=1}^{k}g(Z_{t,j}).$
Because the variables $Z_{t,j}$ are i.i.d. and $\mathcal{G}$ has VC dimension at most $2d$, the Vapnik-Chervonenkis inequality [49, Theorem 2] gives Heoffding-type high probability uniform deviation bounds. Recalling that for all $g\in\mathcal{G}$ we have $\mathbb{E}_{\mu}[g]\leq\epsilon$, we can use relative VC bounds to better control the tail deviations. For instance, [20, Corollary 2] implies that there is a constant $C$ such that with probability at least $1-\delta$,
$\sup_{g\in\mathcal{G}}\sum_{t=1}^{T}\sum_{j=1}^{k}g(Z_{t,j})\leq\epsilon Tk+C\sqrt{\epsilon Tk\left(d\ln\frac{Tk}{d}+\ln\frac{1}{\delta}\right)}+C\ln\frac{Tk}{d}+C\ln\frac{1}{\delta}.$
We now put the two previous estimate with Sections 5.4 and 29, for $k=\frac{1}{\sigma}\ln\frac{T}{\delta}$ and $\epsilon=1/(Tk)$. We note that the bound from Section 5.4 is vacuous if $\frac{1}{\sigma}\ln\frac{T}{\delta}=k\gtrsim T$. Hence, without loss of generality, we can suppose $k\lesssim T$. Similarly, we can suppose that $\ln\frac{T}{\delta}\lesssim\sigma T$. Altogether, this shows that with probability at least $1-2\delta$, we still have
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\lesssim\sqrt{\frac{(d\ln^{2}T+d\ln\ln\frac{1}{\delta}+\ln\frac{1}{\delta})\ln^{3}T}{\sigma}\cdot T}.$ (30)
This ends the proof of Theorem 4.
5.5 Proof of Lemma 13
Fix $p$ and $r$. To prove the desired bound, we first construct a subsequence $(z_{a})_{a}$ of the process $(x_{t})_{t\in[T]}$ that essentially only keeps times for which $\gamma^{(p,r)}(t)\geq q$. For readability, we omit all exponents $(p)$ and $(p,r)$ within this proof from now.
Construction of the alternative smooth process.
Fix a value $q\in[0,1]$ satisfying Eq. 18, and fix the parameter $w\geq 1$. We denote by $(\mathcal{H}_{t})_{t}$ the filtration corresponding to the smooth process $(x_{t})_{t}$. We construct a random subsequence $(z_{a})_{a}$ inductively for $k\in[N_{p}]$. Let $a_{0}=0$. Suppose that for $k\in[N_{p}]$ we have constructed a non-decreasing sequence of indices $a_{0},\ldots,a_{k-1}$ together with elements $z_{1},\ldots,z_{a_{k-1}}$ on $\mathcal{X}$ and values $\gamma_{1},\ldots,\gamma_{a_{k-1}}\in[q,1]$ such that all these random variables are all $\mathcal{H}_{T_{k-1}}$-measurable. We focus on the epoch $E_{k}$ and recall the notation $\mathcal{P}_{k}$ from Eq. 15 for the set of pairs of functions $f,g\in\mathcal{F}$ which had the same values prior $E_{k}$ up to $\epsilon$, as well as the notation $\gamma(t)$ for $t\in E_{k}$ from Eq. 16. We then enumerate
$\{t\in E_{k}:\gamma(t)\geq q\}=:\{t_{1}^{(k)}<\ldots<t_{b_{k}}^{(k)}\}.$
For convenience, for all $l\in[b_{k}]$, we denote $\gamma_{l}^{(k)}:=\gamma(t_{l}^{(k)})$. We then let
$c_{k}:=\min\{b_{k}\}\cup\left\{l\in[b_{k}]:\sum_{l^{\prime}\leq l}\gamma_{l^{\prime}}^{(k)}\geq w\right\}.$ (31)
We next pose $a_{k}=a_{k-1}+c_{k}$ and augment the sequences $z_{1},\ldots,z_{a_{k-1}}$ and $\gamma_{1},\ldots,\gamma_{a_{k-1}}$ as follows
$\left(z_{a_{k-1}+l},\gamma_{a_{k-1}+l}\right):=\left(x_{t_{l}^{(k)}},\gamma_{l}^{(k)}\right),\quad l\in[c_{k}].$
This concludes the construction of the sequence on epoch $k$. We can easily check that all these added random variables are $\mathcal{H}_{T_{k}}$-measurable, which ends the construction of the sequences $(a_{k})_{k\in[N_{p}]}$, $(z_{a})_{a\in[a_{N_{p}}]}$, and $(\gamma_{a})_{a\in[a_{N_{p}}]}$. For convenience, let us denote $A:=a_{N_{P}}$ the random length of these sequences. Note that all constructed quantities $(\gamma_{a})_{a}$ are at least $q$ by construction. Also, for any $a_{k-1}<a\leq a_{k}$, since we added the element $z_{a}=x_{t_{a-a_{k-1}}^{(k)}}$, by definition of $c_{k}$ in Eq. 31, we have
$\sum_{s=a_{k-1}+1}^{a-1}\gamma_{s}<w.$ (32)
The next step is to bound the sum of the quantities $\gamma_{a}$ accumulated on this sequence.
Construction of functions $g_{a}$ for $a\in[T]$
Importantly, we can check that the stochastic process $z_{1},\ldots,z_{A}$ can be constructed online. More precisely, this is a sub-sequence of the smoothed process $x_{1},\ldots,x_{T}$ and is adapted to the filtration $(\mathcal{H}_{t})_{t}$ in the following sense. Knowing whether to add $x_{t}$ in the sequence $z_{1},\ldots,z_{A}$ is $\mathcal{H}_{t-1}$-measurable because this only requires constructing $\gamma(l)$ for all $l\leq t$, which is $\mathcal{H}_{t-1}$-measurable. As a result, $z_{1},\ldots,z_{A}$ is also a $\sigma$-smooth stochastic process for the unknown base measure $\mu$, with the only subtlety being that its horizon is also stochastic. Note that because $(z_{a})_{a\in[A]}$ is a subsequence of $(x_{t})_{t\in[T]}$, we always have $A\leq T$. For convenience, we complete the sequence $z_{1},\ldots,z_{T}$ arbitrarily for $t>A$, for instance with independent samples from $\mu$, as long as the complete process $(z_{a})_{a\in[T]}$ remains $\sigma$-smooth with respect to $\mu$.
For any $a\in[T]$, we define a random function $g_{a}$ as follows. If $a>A$, we can simply pose $g_{a}=0$. Note that knowing whether $a\leq A$ can be done in an online process adapted to the filtration $(\mathcal{H}_{t})_{t\in[T]}$ with the same ideas presented above. Provided $a\leq A$, we denote by $k\in[N_{p}]$ the index such that $a_{k-1}<a\leq a_{k}$ and let $l\in[b_{k}]$ such that we used the time $t_{l}^{(k)}$ to construct $z_{a}=x_{t_{l}^{(k)}}$. We recall that knowing whether we are using $t_{l}^{(k)}$ to construct $z_{a}$ is $\mathcal{H}_{t_{l}^{(k)}-1}$-measurable since we only need to know the past history as well as $\gamma(t_{l}^{(k)})$. By construction, we had $\gamma(t_{l}^{(k)})=\gamma_{l}^{(k)}\geq q>0$. Hence we can fix $f_{l}^{(k)}\in\mathcal{P}_{k}$ such that
$\mathbb{E}\left[\left|f_{l}^{(k)}(x_{t_{l}^{(k)}})-f^{\star}(x_{t_{l}^{(k)}})\right|^{r}\mid\mathcal{H}_{t_{l}^{(k)}-1}\right]$ $\geq(1-\zeta)\sup_{f\in\mathcal{P}_{k}}\mathbb{E}\left[\left|f(x_{t_{l}^{(k)}})-f^{\star}(x_{t_{l}^{(k)}})\right|^{r}\mid\mathcal{H}_{t_{l}^{(k)}-1}\right]$
$=(1-\zeta)\gamma_{a},$ (33)
for a fixed value $\zeta>0$. We then pose $g_{a}:=|f_{l}^{(k)}-f^{\star}|^{r}$.
Upper bound on $\sum_{a=1}^{A}\mathbb{E}\left[g_{a}(z_{a})\mid\mathcal{H}_{t(a)-1}\right]$.
By construction, we can ensure that for all $a\in[T]$, provided $a\leq A$, $g_{a}$ is $\mathcal{H}_{t_{l}^{(k)}-1}$-measurable, where we used the same notations as above for which $z_{a}$ was constructed via $z_{a}=x_{t_{l}^{(k)}}$. To avoid confusions, we denote $t(a):=t_{l}^{(k)}$. In particular, $z_{a}=x_{t(a)}\mid\sigma(z_{a^{\prime}},a^{\prime}<a,g_{a})$ is still $\sigma$-smooth since $\sigma(z_{a^{\prime}},a^{\prime}<a,g_{a})\subset\mathcal{H}_{t(a)-1}$. Last, $A$ is a stopping time for the filtration given by the sigma-algebras $\mathcal{H}_{t(a)-1}$. We are now in position to use Lemma 20 to the rescaled functions $g_{a}/4$ which implies that for a given sequence $z_{1}^{\prime},\ldots,z_{T}^{\prime}$ tangent to $z_{1},\ldots,z_{T}$,
$\sum_{a=1}^{A}\mathbb{E}\left[g_{a}(z_{a})\mid\mathcal{H}_{t(a)-1}\right]$ $\leq 12\sqrt{\frac{2A\ln(eA)}{\sigma}\left(\ln(eA)+\frac{1}{4}\sum_{a=1}^{A}\frac{1}{a}\sum_{s=1}^{a-1}\mathbb{E}\left[g_{a}(z_{s}^{\prime})\mid\mathcal{H}_{t(a)-1}\right]\right)}.$ (34)
We now fix $a\in[T]$ such that $a\leq A$. Using the same notations as before, let $k\in[N_{p}]$ such that $a_{k-1}<a\leq a_{k}$ and $l\in[b_{k}]$ such that we constructed $z_{a}$ via $z_{a}=x_{t_{l}^{(k)}}$. Importantly,
$g_{a}(z_{s})\leq(2\epsilon)^{r},\quad\forall s\leq a_{k-1}.$ (35)
Indeed, recall that $g_{a}=(f_{l}^{(k)}-f^{\star})^{2}$ where $f_{l}^{(k)}\in\mathcal{P}_{k}$. By definition of $\mathcal{P}_{k}$, $f_{l}^{(k)}$ and $f^{\star}$ agree on all queries $x_{t}$ for $t\in[T_{k-1}]$ up to $\epsilon$ in absolute value. Next, let $a_{1}<a$. Assuming that $a\leq A$, we let $k_{1}\in[N_{p}]$ and $l_{1}\in[b_{k_{1}}]$ such that we constructed $z_{a_{1}}:=x_{t_{l_{1}}^{(k_{1})}}$. Note that because $a_{1}<a$, we have $(k_{1},l_{1})<_{lex}(k,l)$, where $<_{lex}$ denotes the lexicographical order. Then, we have
$\mathbb{E}_{z_{a_{1}}^{\prime}}\left[g_{a}(z_{a_{1}}^{\prime})\mid\mathcal{H}_{t(a)-1},a\leq A\right]$ $=\mathbb{E}_{x\sim x_{t_{l_{1}}^{(k_{1})}}\mid\mathcal{H}_{t_{l_{1}}^{(k_{1})}-1}}\left[g_{a}(x)\right]$
$=\mathbb{E}_{x\sim x_{t_{l_{1}}^{(k_{1})}}\mid\mathcal{H}_{t_{l_{1}}^{(k_{1})}-1}}\left[\left|f_{l}^{(k)}(x)-f^{\star}(x)\right|^{r}\right]$
$\leq\gamma\left(t_{l_{1}}^{(k_{1})}\right)=\gamma_{a_{1}}.$ (36)
In the last inequality, we used the definition of the function $\gamma(\cdot)$ from Eq. 16 and the fact that by construction $f_{l}^{(k)}\in\mathcal{P}_{k}\subset\mathcal{P}_{k_{1}}$ (note that $\mathcal{P}_{k}$ only has more constraints on the functions compared to $\mathcal{P}_{k_{1}}$) . Putting these equations together, we obtain
$\sum_{s=1}^{a-1}$ $\mathbb{E}\left[g_{a}(z_{s}^{\prime})\mid\mathcal{H}_{t(a)-1},a\leq A\right]\overset{(i)}{\leq}\sum_{s=a_{k-1}+1}^{a-1}\gamma_{s}+\sum_{s=1}^{a_{k-1}}\mathbb{E}\left[g_{a}(z_{s}^{\prime})\mid\mathcal{H}_{t(a)-1},a\leq A\right]$
$\overset{(ii)}{\leq}w+\underbrace{\mathbb{E}_{z_{1}^{\prime},\ldots,z_{a_{k-1}}^{\prime}}\left[\sum_{s=1}^{a_{k-1}}g_{a}(z_{s}^{\prime})-2g_{a}(z_{s})\mid\mathcal{H}_{t(a)-1},a\leq A\right]}_{E(a)}+(2\epsilon)^{r}a_{k-1}.$ (37)
In $(i)$ we used Eq. 36 and in $(ii)$ we used Eqs. 35 and 32. Now note that conditionally on $a\leq A$ and $\mathcal{H}_{t(a)-1}$, we have
$\sum_{s=1}^{a_{k-1}}g_{a}(z_{s}^{\prime})-2g_{a}(z_{s})$ $=\sum_{s=1}^{a_{k-1}}\left|f_{l}^{(k)}(z_{s}^{\prime})-f^{\star}(z_{s}^{\prime})\right|^{r}-2\left|f_{l}^{(k)}(z_{s})-f^{\star}(z_{s})\right|^{r}$
$\leq\sup_{f\in\mathcal{F}}\sum_{s=1}^{a_{k-1}}|f(z_{s}^{\prime})-f^{\star}(z_{s}^{\prime})|^{r}-2|f(z_{s})-f^{\star}(z_{s})|^{r}$
$\overset{(i)}{\leq}\sup_{\tilde{a}\leq T}\sup_{f\in\mathcal{F}}\sum_{s=1}^{\tilde{a}}|f(z_{s}^{\prime})-f^{\star}(z_{s}^{\prime})|^{r}-2|f(z_{s})-f^{\star}(z_{s})|^{r}.$
Note that we perform step $(i)$ because $a_{k-1}$ is not a fixed horizon a priori: it may depend on the elements of the smooth sequence $z_{b}$ for $b>a_{k-1}$ (precisely, the elements $a_{k-1}<b\leq a$). We now bound the right-hand side using a high-probability variant of [15, Theorem 2], given in Lemma 22. Precisely, we apply Lemma 22 to the function class $\mathcal{F}_{p}:=\{\frac{1}{r}|f-f^{\star}|^{r}:f\in\mathcal{F}\}$ with the parameter $c=1/2$. Together with the union bound this implies that with probability at least $1-\delta^{2}$,
$\sup_{\tilde{a}\leq T}\sup_{f,h\in\mathcal{F}}\sum_{s=1}^{\tilde{a}}|f(z_{s}^{\prime})-h(z_{s}^{\prime})|^{r}-2|f(z_{s})-h(z_{s})|^{r}\leq rC_{1}\left(\ln\mathbb{E}_{\mu}\left[W_{4T\ln(\frac{T}{\delta})/\sigma}\left(\frac{1}{3}\mathcal{F}_{p}\right)\right]+\ln\frac{T}{\delta}\right),$ (38)
for some universal constant $C_{1}\geq 1$. Here we used the fact that the Wills functional $W_{m}(\mathcal{F}_{p})$ is non-decreasing in $m$ (e.g. see [15, Lemma 10]) and that $a\leq T$. Now note that the function $\psi:z\in[-1,1]\mapsto\frac{1}{r}|x|^{r}$ is $1$-Lipschitz. Hence [41, Theorem 4.1] implies that $W_{m}(\frac{1}{3}\mathcal{F}_{p})\leq W_{m}(\tilde{\mathcal{F}})$, where ~=$\tilde{\mathcal{F}}=\{f-f^{\star}:f\in\mathcal{F}\}$. Next, because the Wills functional is invariant under translation from [41, Proposition 3.1.5], we finally obtain
$W_{m}\left(\frac{1}{3}\mathcal{F}_{p}\right)\leq W_{m}(\mathcal{F}),\quad m\geq 1.$
Denote by $\mathcal{E}_{\delta^{2}}$ the event when Eq. 38 holds. Then,
$E(a)\leq rC_{1}\left(\ln\mathbb{E}_{\mu}\left[W_{4T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+2\ln\frac{T}{\delta}\right)+T\mathbb{P}\left(\mathcal{E}_{\delta^{2}}^{c}\mid\mathcal{H}_{t(a)-1},a\leq A\right).$
where the probability on the last term is taken over $z_{1}^{\prime},\ldots,z_{T}^{\prime}$. Now by Markov’s inequality, we have
$\mathbb{P}_{z_{1},\ldots,z_{T}}\left[\mathbb{P}_{z_{1}^{\prime},\ldots,z_{T}^{\prime}}\left(\mathcal{E}_{\delta^{2}}^{c}\mid\mathcal{H}_{t(a)-1},a\leq A\right)\geq\delta\right]\leq\frac{\mathbb{P}(\mathcal{E}_{\delta^{2}}^{c})}{\delta}\leq\delta,$
Denote by $\mathcal{F}_{\delta}(a)$ the complementary event, which has probability at least $1-\delta$. On this event, the previous bound from Eq. 37 implies that
$\sum_{s=1}^{a-1}\mathbb{E}\left[g_{a}(z_{s}^{\prime})\mid\mathcal{H}_{t(a)-1},a\leq A\right]\leq w+rC_{1}\left(\ln\mathbb{E}_{\mu}\left[W_{4T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}\right)+\delta T+(2\epsilon)^{r}a.$
Plugging this bound into Eq. 34 shows that on $\bigcap_{a\leq T}\mathcal{F}_{\delta}(a)$ which has probability at least $1-\delta T$,
$\sum_{a=1}^{A}\mathbb{E}\left[g_{a}(z_{a})\mid\mathcal{H}_{t(a)-1}\right]$
$\leq 12\left(\frac{2A\ln(eA)}{\sigma}\left[\ln(eA)+\left(rC_{1}\ln\mathbb{E}_{\mu}\left[W_{4T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+rC_{1}\ln\frac{T}{\delta}+w+\delta T\right)\sum_{a=1}^{T}\frac{1}{4a}\right]\right.$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{(2\epsilon)^{r}A^{2}\ln(eA)}{2\sigma}\right)^{1/2}$
$\leq C\ln T\sqrt{\frac{rA}{\sigma}\left(\ln\mathbb{E}_{\mu}\left[W_{4T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w+\delta T\right)}+6A\sqrt{(2\epsilon)^{r}\frac{2\ln(eT)}{\sigma}},$ (39)
for some universal constant $C\geq 1$. In the last inequality, we used $A\leq T$ and the inequality $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for all $a,b\geq 0$. For convenience, we introduce the notation
$C_{w,\delta}(T):=C\ln T\sqrt{\frac{r}{\sigma}\left(\ln\mathbb{E}_{\mu}\left[W_{4T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w+\delta T\right)}.$
Lower bound on $\sum_{a=1}^{A}\mathbb{E}\left[g_{a}(z_{a})\mid\mathcal{H}_{t(a)-1}\right]$.
We now turn to the lower bound. Note that for any $a\in[T]$ provided that $a\leq A$, using the same notations as above we have
$\mathbb{E}\left[g_{a}(z_{a})\mid\mathcal{H}_{t(a)-1},a\leq A\right]=\mathbb{E}\left[\left|f_{l}^{(k)}(x_{t_{l}^{(k)}})-h_{l}^{(k)}(x_{t_{l}^{(k)}})\right|^{r}\mid\mathcal{H}_{t(a)-1},a\leq A\right]\overset{(i)}{\geq}(1-\zeta)\gamma_{a},$ (40)
where in $(i)$ we used Eq. 33. As a result,
$\sum_{a=1}^{A}\mathbb{E}\left[g_{a}(z_{a})\mid\mathcal{H}_{t(a)-1}\right]\geq(1-\zeta)\sum_{a=1}^{A}\gamma_{a}.$
Putting the two bounds together.
Putting together this lower bound with the upper bound from Eq. 39, we obtain that with probability at least $1-\delta$
$(1-\zeta)\sum_{a=1}^{A}\gamma_{a}$ $\leq C_{w,\delta/T}(T)\sqrt{A}+6A\sqrt{(2\epsilon)^{r}\frac{2\ln(eT)}{\sigma}}$
$\overset{(i)}{\leq}C_{w,\delta/T}(T)\sqrt{\frac{1}{q}\sum_{a=1}^{A}\gamma_{a}}+\frac{6}{q}\sqrt{(2\epsilon)^{r}\frac{2\ln(eT)}{\sigma}}\sum_{a=1}^{A}\gamma_{a}$
$\overset{(ii)}{\leq}C_{w,\delta/T}(T)\sqrt{\frac{1}{q}\sum_{a=1}^{A}\gamma_{a}}+\frac{1}{2}\sum_{a=1}^{A}\gamma_{a}$
where in $(i)$ we recalled that for all $a\leq A$, we have $p_{a}\geq q$ and in $(ii)$ we used the assumption on $q$ from Eq. 18. This holds for any $\zeta>0$, which implies that there exists a universal constant $C_{1}\geq 1$ such that for any $\delta\in(0,1/2]$, with probability at least $1-\delta$,
$\sum_{a=1}^{A}\gamma_{a}\leq\frac{2C^{2}_{w,\delta/T}(T)}{q}\leq\frac{C_{1}r\ln^{2}T}{q\sigma}\left(\ln\mathbb{E}_{\mu}\left[W_{8T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w\right).$
Going back to the construction of the sequence $z_{1},\ldots,z_{A}$, for any epoch $E_{k}^{(p)}$, in its construction we always try to include as many times $t_{1}^{(k)},t_{2}^{(k)},\ldots$ as possible until the threshold $w$ for the sum of their probabilities $\gamma_{l}^{(k)}$ is passed. Denote by $\mathcal{K}\subset[N_{p}]$ the set of all epochs $k$ for which not all elements $t_{l}^{(k)}$ for $l\in[b_{k}]$ have been used, that is $\mathcal{K}=\{k\in[N_{p}]:c_{k}<b_{k}\}$. On one hand, if $k\notin\mathcal{K}$, Eq. 32 implies that
$\sum_{t\in E_{k}}\gamma(t)\mathbbm{1}_{\gamma(t)\geq q}=\sum_{l\in[b_{k}]}\gamma_{l}^{(k)}\leq w+1\leq 2w.$
In the last inequality we used $\gamma_{b_{k}}^{(k)}\leq 1$. On the other hand, for any $k\in\mathcal{K}$,
$\sum_{a=a_{k-1}+1}^{a_{k}}\gamma_{a}>w.$
Therefore, with probability at least $1-\delta$,
$|\mathcal{K}|<\frac{1}{w}\sum_{k\in[N_{p}]}\sum_{a=a_{k-1}+1}^{a_{k}}\gamma_{a}$ $=\frac{1}{w}\sum_{a=1}^{A}\gamma_{a}$
$\leq\frac{C_{1}r\ln^{2}T}{q\sigma w}\left(\ln\mathbb{E}_{\mu}\left[W_{8T\ln(\frac{T}{\delta})/\sigma}\left(\mathcal{F}\right)\right]+\ln\frac{T}{\delta}+w\right).$
Considering $2w$ instead of $w$ and up to changing the constant $C_{1}$, this ends the proof of the first claim.
To prove the second claim, let $w\geq 2$ and $q\in(0,1]$ satisfying Eq. 19. For any $k\in[N_{p}]$, we have
$\Gamma_{k}=\sum_{t\in E_{k}}\gamma(t)\leq\frac{q}{2}(T_{k}-T_{k-1})+\sum_{t\in E_{k}}\gamma(t)\mathbbm{1}_{\gamma(t)\geq q/2}.$
Applying the bound proved above for $q/2$ together with the fact that $w\leq\frac{q}{2}(L_{p}-1)\leq\frac{q}{2}(T_{k}-T_{k-1})$ ends the proof of the second claim.
Acknowledgments
The author is very grateful to Abhishek Shetty and Adam Block for bringing this problem to his attention and for insightful discussions. This work was supported by a Columbia Data Science Institute postdoctoral fellowship.
Appendix A Bounds on the Wills functional
We recall the definition of the Wills functional
$W_{m,Z}(\mathcal{F}):=\mathbb{E}_{\xi}\left[\exp\left(\sup_{f\in\mathcal{F}}\sum_{i=1}^{m}\xi_{i}f(Z_{i})-\frac{1}{2}f^{2}(Z_{i})\right)\right],$
where $\xi$ is a vector of $m$ i.i.d. standard Gaussians. A first way to bound the Wills functional is to bound either the Gaussian complexity or the Rademacher complexity of the function class. We recall their definitions below.
$\mathcal{R}_{m,Z}(\mathcal{F})$ $:=\mathbb{E}_{\epsilon}\left[\sup_{f\in\mathcal{F}}\sum_{i=1}^{m}\epsilon_{i}f(Z_{i})\right]$
$\mathcal{G}_{m,Z}(\mathcal{F})$ $:=\mathbb{E}_{\xi}\left[\sup_{f\in\mathcal{F}}\sum_{i=1}^{m}\xi_{i}f(Z_{i})\right],$
where $\xi$ is a vector of $m$ i.i.d. standard Gaussians and $\epsilon$ is a vector of $m$ i.i.d. Rademacher variables. We may omit the dependency in the values $Z=(Z_{1},\ldots,Z_{m})$ when clear from context. We have the following
Proposition 16 (Proposition 3.2 of [41], Proposition 3 of [15], Exercise 5.5 of [52]).
For any function class $\mathcal{F}$, $m\in\mathbb{N}$, and values $Z_{1},\ldots,Z_{m}\in\mathcal{X}$, we have
$\ln W_{m}(\mathcal{F})\leq\mathcal{G}_{m}(\mathcal{F})\lesssim\sqrt{\ln m}\cdot\mathcal{R}_{m}(\mathcal{F}).$
More precisely, [41] gave a characterization for the Wills functional, in terms of the local Gaussian complexity and covering numbers which we now define. Having fixed $Z_{1},\ldots,Z_{m}$, we introduce the notation $\mu_{m}=\frac{1}{m}\sum_{i=1}^{m}\delta_{Z_{i}}$ for the uniform distribution on the values $Z_{1},\ldots,Z_{n}$ and define the norm $\|f\|_{L_{2}(\mu_{m})}:=(\mathbb{E}_{Z\sim\mu_{m}}|f(Z)|^{2})^{1/2}$ for any function $f$. The local Gaussian complexity is defined as follows
$\mathcal{G}_{m,Z}(\mathcal{F},r):=\sup_{f_{0}\in\mathcal{F}}\mathcal{G}_{m,Z}(B_{r}(f_{0};\mathcal{F})),$
where $B_{r}(f_{0};\mathcal{F})=\{f\in\mathcal{F}:\|f-g\|_{L_{2}(\mu_{m})}\leq r\}$ is the ball within $\mathcal{F}$ centered at $f_{0}$ of radius $r$. Again, we may omit the dependency in $Z$. The covering number $\mathcal{N}_{2,m}(\mathcal{F},r)$ is defined as the minimal cardinality of an $r$-cover of $\mathcal{F}$ with respect to $\|\cdot\|_{L_{2}(\mu_{m})}$. As a remark, these notations differ from those in Theorem 17 by a factor $\sqrt{m}$ for the scale $r$. This choice of scaling will be easier to work with when computing covering numbers.
Theorem 17 (Theorem 4.2 of [41]).
There exist constants $c,C>0$ such that the following holds. For any function class $\mathcal{F}$, $m\in\mathbb{N}$, and values $Z_{1},\ldots,Z_{m}\in\mathcal{X}$, we have
$c\cdot\inf_{r>0}\left\{\mathcal{G}_{m}(\mathcal{F},r)+\ln\mathcal{N}_{2,m}(\mathcal{F},r)\right\}\leq\ln W_{m}(\mathcal{F})\leq C\cdot\inf_{r>0}\left\{\mathcal{G}_{m}(\mathcal{F},r)+\ln\mathcal{N}_{2,m}(\mathcal{F},r)\right\}.$
In particular, we obtain the following bounds for classical behaviors of function classes.
Proposition 18.
Fix any values $Z_{1},\ldots,Z_{m}\in\mathcal{X}$. If $\mathcal{F}$ is finite, then $\ln W_{m}(\mathcal{F})\lesssim\ln|\mathcal{F}|$. If $\mathcal{F}$ has finite VC dimension $d$, then $\ln W_{m}(\mathcal{F})\lesssim d\ln m$.
More generally, for any $r>0$,
$\mathcal{G}_{m}(\mathcal{F},r)\lesssim\inf_{0\leq r^{\prime}\leq r}\left\{r^{\prime}m+\sqrt{m}\cdot\int_{r^{\prime}}^{r}\sqrt{\mathsf{fat}_{\mathcal{F}}(\epsilon)}\ln\frac{16\cdot\mathsf{fat}_{\mathcal{F}}(\epsilon)}{\epsilon}d\epsilon\right\}$ (41)
In particular, if there exists $d\geq 1$ such that for all $r>0$, one has $\mathsf{fat}_{\mathcal{F}}(r)\leq d\ln\frac{1}{r}$, we have
$\ln W_{m}(\mathcal{F})\lesssim d\ln^{3}(dm).$
In particular, if there exists some $\gamma>1$ such that for any $r>0$, $\mathsf{fat}_{\mathcal{F}}(r)\leq\gamma r^{-p}$, for all $r>0$,
$\ln W_{m}(\mathcal{F})\lesssim_{p}\begin{cases}\gamma^{\frac{2}{2+p}}m^{\frac{p}{2+p}}\cdot\ln^{\frac{4}{2+p}}(\gamma m)&0<p<2\\ \sqrt{\gamma m}\cdot\ln^{2}(\gamma m)+\gamma\ln^{2}\gamma&p=2\\ \gamma^{\frac{1}{p}}m^{1-\frac{1}{p}}\cdot\ln^{\frac{2}{p}}(\gamma m)+\gamma\ln^{2}\gamma&p>2.\end{cases}$
where $\lesssim_{p}$ only hides factors that depend (possibly exponentially) only on $p$. These bounds can be simplified as follows
$\ln W_{m}(\mathcal{F})\lesssim_{\gamma,p}m^{\alpha(p)}\ln^{2}m,\quad\text{where}\quad\alpha(p):=:=\begin{cases}\frac{p}{2+p}&0<p\leq 2\\ 1-\frac{1}{p}&p\geq 2,\end{cases}$
where $\lesssim_{p,\gamma}$ hides factors and additive terms depending on $p,\gamma$ only.
Proof  For function classes $\mathcal{F}$ with finite VC dimension $d$, Sauer-Shelah’s Lemma 1 implies that $\ln\mathcal{N}_{2}(\mathcal{F},r)\lesssim d\ln m$ for any $r\in[0,1]$, which directly implies that $\ln W_{m}(\mathcal{F})\lesssim d\ln m$. Similarly, for any finite class $\mathcal{F}$, we obtain $\ln W_{m}(\mathcal{F})\lesssim\ln|\mathcal{F}|$.
Next, from [40, Theorem 3.2], we have for any $r>0$,
$\ln\mathcal{N}_{2,m}(\mathcal{F},r)\lesssim\mathsf{fat}_{\mathcal{F}}(r/8)\ln^{2}\left(\frac{2\mathsf{fat}_{\mathcal{F}}(r/8)}{r}\right).$ (42)
We can combine this estimate with the chaining bounds for Gaussian complexities from [40, Lemma 3.7] together with the fact $\mathcal{N}_{2,m}(B_{r}(f_{0};\mathcal{F}),r^{\prime})=1$ for all $r^{\prime}>r$ and $f_{0}\in\mathcal{F}$, which implies the desired bound on the local Gaussian complexity Eq. 41.
Suppose that we have $\mathsf{fat}_{\mathcal{F}}(r)\leq d\ln\frac{1}{r}$ for all $r>0$. Then, we can choose $r=1/m$ in Theorem 17 and $r^{\prime}=r$ in Eq. 41 which gives the desired result.
Now suppose that $\mathsf{fat}_{\mathcal{F}}(r)\leq\gamma r^{-p}$ for all $r>0$ for some $\gamma>1$ and $p>0$. Then, for $r\in(0,1]$, Eq. 41 yields
$\mathcal{G}_{m}(\mathcal{F},r)\lesssim_{p}\begin{cases}\sqrt{\gamma m}\cdot r^{1-\frac{p}{2}}\ln\frac{8\gamma}{r}&0<p<2\\ \min\left\{rm,\sqrt{\gamma m}\cdot\ln m\cdot\ln(\gamma m)\right\}&p=2\\ \min\left\{rm,\gamma^{\frac{1}{p}}m^{1-\frac{1}{p}}\cdot\ln^{\frac{2}{p}}(\gamma m)\right\}&p>2.\end{cases}$
This can be obtained directly from Eq. 41 by plugging in the value $r^{\prime}=0$ for $20<p<$ 2. For $p=2$, we take $r^{\prime}=\min\left\{r,\sqrt{\gamma/m}\right\}$. Last, for $p>2$, we take $r^{\prime}=\min\left\{r,(\gamma\ln^{2}(\gamma m)/m)^{\frac{1}{p}}\right\}$.
We then use Theorem 17 together with Eq. 42 and the previous estimates on the local Gaussian complexity to obtain the desired bound on the Wills functional $W_{m}(\mathcal{F})$. For $20<p<$ 2, we used the value $r=\left(\gamma\ln^{2}(\gamma m)/m\right)^{\frac{1}{p+2}}$. For $p\geq 2$, we used the value $r=1$. $\blacksquare$
Appendix B Learning with expert advice guarantee for A-Exp
In this section, we prove Lemma 7. Note that A-Exp proceeds by periods $k\geq 1$. Let $T_{0}=0$ and denote by $T_{k}$ the end of period $k$ for $k\geq 1$. That is,
$T_{k}=\min\left\{t>T_{k-1}:\sum_{l=T_{k-1}+1}^{t}\sum_{i\in[K]}p_{l,i}r_{l,i}^{2}>\Delta_{max,k}=4^{k-1}\right\},\quad k\geq 1$
On period $(T_{k-1},T_{k}]$, A-Exp exactly implements the exponentially weighted forecaster with parameter $\eta_{k}=\sqrt{2\ln K/(\Delta_{max,k}+1)}$. Hence, we can use Eq. 1 to bound the regret accumulated on this period which gives for all $T\in(T_{k-1},T_{k}]$,
$\sum_{t=T_{k-1}+1}^{T}\mathbb{E}_{\hat{i}_{t}}[\ell_{t,\hat{i}_{t}}\mid\mathcal{H}_{t}]-\min_{i\in[K]}\sum_{t=T_{k-1}+1}^{T}\ell_{t,i}$ $\leq\frac{\ln K}{\eta_{k}}+\frac{\eta_{k}}{2}\sum_{t=T_{k-1}+1}^{T}\sum_{i\in[K]}p_{t,i}r_{t,i}^{2}$
$\overset{(i)}{\leq}\frac{\ln K}{\eta_{k}}+\frac{\eta_{k}}{2}(\Delta_{max,k}+1)$
$=\sqrt{2(\Delta_{max,k}+1)\ln K}=\sqrt{2(4^{k-1}+1)\ln K}\leq 2^{k}\sqrt{\ln K},$
where in $(i)$ we used the fact that $r_{T_{k},i}\in[0,1]$ for all $i\in[K]$. Now for $T\geq 1$ denote by $k$ the last period, such that $T\in(T_{k-1},T_{k}]$. Provided $k\geq 2$, we can sum the previous equations for periods $k^{\prime}\leq k$ to obtain
$PReg(T)\leq\sum_{k^{\prime}\leq k}2^{k}\sqrt{\ln K}\leq 2^{k+1}\sqrt{\ln K}\overset{(ii)}{\leq}8\sqrt{\Delta_{T}\ln K}.$
In $(ii)$ we used the fact that if $k\geq 2$ then
$\Delta_{T}\geq\sum_{t=T_{k-2}+1}^{T_{k-1}}\sum_{i\in[K]}p_{t,i}r_{t,i}^{2}>\Delta_{max,k-1}=4^{k-2}.$
If $k=1$, we have directly $PReg(T)\leq 2\sqrt{\ln K}$. This ends the proof for the bound on the pseudo-regret.
To obtain high-probability bounds on the regret $Reg(T)$, we could simply use Azuma-Hoeffding’s inequality. However, this would add an additional term $\sqrt{T\ln\frac{1}{\delta}}$ that is prohibitive for our bounds: potentially we have $\Delta_{T}\ll T$. Instead, we use Freedman’s inequality which gives a more precise control on tail probabilities for martingales. Lemma 19 applied with $Z_{t}=r_{t,\hat{i}_{t}}^{2}-\mathbb{E}_{\hat{i}_{t}}[r_{t,\hat{i}_{t}}^{2}\mid\mathcal{H}_{t}]$ for $t\in[T]$ and $\eta=1/2$ implies that with probability at least $1-\delta$,
$Reg(T)\overset{(i)}{\leq}PReg(T)+\frac{1}{2}\sum_{t=1}^{T}\mathbb{E}[r_{t,\hat{i}_{t}}^{4}\mid\mathcal{H}_{t}]+2\ln\frac{1}{\delta}\overset{(ii)}{\leq}\frac{3}{2}PReg(T)+2\ln\frac{1}{\delta},$
where in $(i)$ we used the fact that $Var(Y)\leq\mathbb{E}[Y^{2}]$ for any random variable $Y$ and in $(ii)$ we used the fact that $|r_{t,i}|\leq 1$ for all $i\in[K]$ and $t\in[T]$. This ends the proof.
Appendix C Proof of the regret lower bound
In this section, we prove that the regret bound from Theorem 4 is tight up to logarithmic terms. We recall the statement of the lower bound below.
See 5
Proof  The template function class that we use are simply the threshold functions on $[0,1]\mapsto\{0,1\}$, which have VC dimension one. To extend this to a function class with VC dimension $d$, we take $d$ copies. That is, we pose $\mathcal{X}=\{1,\ldots,d\}\times[0,1]=[d]\times[0,1]$ and we let
:=$\mathcal{F}:=\left\{f_{\boldsymbol{\theta}}:(k,x)\in\mathcal{X}\mapsto\mathbbm{1}[x\geq\theta_{k}],\;\boldsymbol{\theta}\in[0,1]^{d}\right\}.$
For convenience, we define $\bar{x}=(1,0)$, where the value $1$ was chosen arbitrarily, we also let $\mathcal{X}_{k}=\{k\}\times[0,1]$. By definition, we have $\mathcal{X}=\mathcal{X}_{1}\sqcup\ldots\sqcup\mathcal{X}_{d}$.
Now fix a horizon $T\geq 1$ and $\sigma\in(0,1)$. Suppose for now that
$T>\frac{4d(1-\sigma)}{\sigma}.$ (43)
We now fix a parameter $q=\sqrt{\frac{d(1-\sigma)}{\sigma T}}$ and let $N=\left\lfloor qT/d\right\rfloor$. Note that from the assumption on $T$, we have $q<1/2$. Next, suppose that $N\leq 1$. Then, this corresponds to $q\leq 2d/T$. Classical lower bounds for VC classes show that we can construct a distribution $\mu$ (uniform on $d$ shattered points), which corresponds to $\sigma=1$ together with a function $f^{\star}\in\mathcal{F}$ such that with $x_{t}\overset{iid}{\sim}\mu$, the expected number of mistakes of any algorithm is at least $\min(d,T)/4$. Now because $qT\leq 2d$, this directly implies the desired result when $N\leq 1$.
From now, we suppose that $N\geq 2$. Let $\boldsymbol{\epsilon}=(\epsilon_{k,t})_{k\in[d],t\in[N]}$ be a sequence of i.i.d. uniform variables on $\{0,1\}$. We now construct a generating process for the sequence $(x_{t},y_{t})_{t\in[T]}$ coupled with $\boldsymbol{\epsilon}$. In addition to the variables $(x_{t},y_{t})_{t\in[T]}$, the process also iteratively constructs variables $a_{k,t}<b_{k,t}$ for $k\in[d]$ and $t\in[T]$. For each $k\in[d]$, the interval $\{k\}\times(a_{k,t},b_{k,t})$ will intuitively represent the region of $\mathcal{X}_{k}$ on which the learner does not have information yet at the beginning of round $t$.
We initialize the process at time $t=0$ by setting $a_{k,1}=0$ and $b_{k,1}=1$ for all $k\in[d]$. We also initialize index variables $i(k,1)=1$ for all $k\in[d]$. Suppose that we have constructed $a_{k,t},b_{k,t}$ for $k\in[d]$ at some iteration $t\in[T]$, as well as the indices $i(k,t)$ for $k\in[d]$. We then define the distribution
$\mu_{t}:=(1-q)\delta_{\bar{x}}+\sum_{k=1}^{d}\frac{q}{d}\left(\delta_{(k,(a_{k,t}+b_{k,t})/2)}\mathbbm{1}_{i(k,t)\leq N}+\delta_{\bar{x}}\mathbbm{1}_{i(k,t)>N}\right),$ (44)
where $\delta_{z}$ denotes the Dirac distribution at $z$, and $q\in(0,1)$ is a fixed probability value. We then sample $x_{t}\sim\mu_{t}$ independently from $\boldsymbol{\epsilon}$ and let
$y_{t}:=\begin{cases}0&\text{if }x_{t}=\bar{x}\\ \epsilon_{i(k,t)}&x_{t}\in\mathcal{X}_{k}.\end{cases}$
We then pose for all $k\in[d]$,
$(a_{k,t+1},b_{k,t+1}):=\begin{cases}(a_{k,t},b_{k,t})&\text{if }x_{t}=\bar{x}\text{ or }x_{t}\notin\mathcal{X}_{k}\\ (a_{k,t},(a_{k,t}+b_{k,t})/2)&\text{if }x_{t}=(k,(a_{k,t}+b_{k,t})/2)\text{ and }\epsilon_{i(k,t)}=1\\ ((a_{k,t}+b_{k,t})/2,b_{k,t})&\text{if }x_{t}=(k,(a_{k,t}+b_{k,t})/2)\text{ and }\epsilon_{i(k,t)}=0.\end{cases}$
Last, we pose for all $k\in[d]$,
$i(k,t+1):=\begin{cases}i(k,t)&\text{if }x_{t}=\bar{x}\text{ or }x_{t}\notin\mathcal{X}_{k}\\ i(k,t)+1&\text{otherwise}.\end{cases}$
This concludes the construction of the process $(x_{t},y_{t})_{t\in[T]}$. Note that by construction, whenever $x_{t}\neq\bar{x}_{t}$, a fresh random variable from $\boldsymbol{\epsilon}$ is used to define $y_{t}$. In particular, we can check that conditionally on the history up to time $t$, we have $y_{t}=0$ if $x_{t}=\bar{x}$ and $y_{t}\sim\mathcal{U}(\{0,1\})$ otherwise. In particular, we always have
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}_{\hat{y}_{t}\neq y_{t}}\right]$ $=\mathbb{E}\left[\sum_{t=1}^{T}\mathbb{E}\left[\mathbbm{1}_{\hat{y}_{t}\neq y_{t}}\mid\hat{y}_{t},\;(x_{l},y_{l})_{l\leq t-1}\right]\right]$
$\geq\frac{1}{2}\mathbb{E}\left[\sum_{t=1}^{T}\mathbb{E}\left[\mathbbm{1}_{x_{t}\neq\bar{x}}\mid(x_{l},y_{l})_{l\leq t-1}\right]\right]$
$=\frac{1}{2}\mathbb{E}\left[\sum_{t=1}^{T}\frac{q}{d}\sum_{k\in[d]}\mathbbm{1}_{i(k,t)\leq N}\right]$
=$\overset{(i)}{=}\frac{q}{2}\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}_{i(1,t)\leq N}\right]=\frac{q}{2}\mathbb{E}_{Z\sim\text{NB}(N,q/d)}\left[\max(N+Z,T)\right],$
where $\text{NB}(r,p)$ denotes the negative binomial distribution with $r$ successes and probability of success $p$. Indeed, $i(1,t)$ grows when $x_{t}=(1,(a_{1,t}+b_{1,t})/2)$, which has probability $q/d$ conditionally on the history. In $(i)$ we use the fact that all coordinates are treated symmetrically. From now let $Z$ be a random variable distributed according to $\text{NB}(N,q/d)$. From [51] since $\mathbb{E}[Z]=\frac{N(1-q/d)}{q/d}>N$, letting $\eta$ be the median of $Z$, we have
$T-N\geq\mathbb{E}[Z]\geq\eta\geq 1+\frac{N-1}{N}\mathbb{E}[Z]=2+\frac{(N-1)d}{q}-N.$
As a result, we have
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}_{\hat{y}_{t}\neq y_{t}}\right]\geq\frac{q(N+\eta-1)}{4}\geq\frac{(N-1)d}{4}.$
By the law of total probabilities, there is a realization of $\boldsymbol{\epsilon}$ which we denote $\tilde{\boldsymbol{\epsilon}}$ such that
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}_{\hat{y}_{t}\neq y_{t}}\mid\boldsymbol{\epsilon}=\tilde{\boldsymbol{\epsilon}}\right]\geq\frac{(N-1)d}{4}.$
By construction of the process, to each realization of $\boldsymbol{\epsilon}$ is associated a function in class $f_{\boldsymbol{\theta}(\boldsymbol{\epsilon})}\in\mathcal{F}$ that realizes all the values $(x_{t},y_{t})$. Indeed, we can take for instance
$\theta(\boldsymbol{\epsilon})_{i}:=\frac{1}{2^{T+1}}+\sum_{t=1}^{T}\frac{1-\epsilon_{k,t}}{2^{k}}.$
Indeed, the main point is that defining the intervals $[a_{k,t},b_{k,t}]$ for $t\in[T]$, only used the variables $\epsilon_{k,t}$ for $t\in[T]$. These are only updated when we sample $x_{t}=(k,(a_{k,t}+b_{k,t})/2)$ in which case we use the first value within $\{\tilde{\epsilon}_{k,1},\ldots,\tilde{\epsilon}_{k,T}\}$ that was not used up to this point. In particular, this implies that the number of possible values that the sequence $(x_{t})_{t\in[T]}$ can take is at most $1+dN$ where the term $1$ corresponds to the value $\bar{x}$. For convenience, let $\nu$ denote the uniform distribution on these $dN$ points where we deleted the value $\bar{x}$.
It now remains to argue that the sequence $(x_{t})_{t\in[T]}$ constructed with $\tilde{\boldsymbol{\epsilon}}$ is $\sigma$-smooth. To do so, we construct the measure
$\mu:=\sigma\delta_{\bar{x}}+(1-\sigma)\nu.$
Importantly, this distribution is fixed a priori (it does not depend on the actions of the learner, only on $\tilde{\boldsymbol{\epsilon}}$ that is fixed). Given its definition in Eq. 44, to check that at any time $t\in[T]$, the distribution $\mu_{t}$ is $\sigma$-smooth compared to the base measure $\mu$, it suffices to check that
$\frac{q/d}{(1-\sigma)/(dN)}=\frac{qN}{1-\sigma}\leq\frac{q^{2}T}{d(1-\sigma)}\leq\frac{1}{\sigma}.$
In the last inequality we used the definition of $q$. As a summary, the sequence $(x_{t})_{t\in[T]}$ is $\sigma$-smooth compared to $\mu$ and using the realizable values $y_{t}=f_{\boldsymbol{\theta}(\tilde{\boldsymbol{\epsilon}})}(x_{t})$, we obtained
$\mathbb{E}\left[\sum_{t=1}^{T}\mathbbm{1}_{\hat{y}_{t}\neq y_{t}}\right]\geq\frac{(N-1)d}{4}\geq\frac{qT}{12}=\frac{1}{12}\sqrt{\frac{dT(1-\sigma)}{\sigma}}.$
In the second inequality we used the assumption that $N\geq 2$ to show that $N-1\geq qT/3d$.
We now consider the case when Eq. 43 is not necessarily satisfied. Then, with $T_{0}=\left\lceil\frac{4d(1-\sigma)}{\sigma}\right\rceil$, the previous arguments imply that for some realizable data and a $\sigma$-smooth adversary, we have
$\mathbb{E}\left[\sum_{t=1}^{T_{0}}\mathbbm{1}_{\hat{y}_{t}\neq y_{t}}\right]\geq\frac{1}{12}\sqrt{\frac{dT_{0}(1-\sigma)}{\sigma}}\geq\frac{T_{0}}{24}.$
As a result, considering the interval of time that incurred the most regret, this shows that for all $T\leq T_{0}$, there is a $\sigma$-smooth realizable adversary under which the expected number of mistakes for any learning algorithm is at least $T/24$. This ends the proof. $\blacksquare$
Appendix D Proofs from Section 4
In this section, we prove the results related to the simplified algorithm Cover. These are essentially simplified proofs of their counterparts for the main proofs from Section 5, hence we will only highlight the main differences.
Proof of Proposition 9  Lemma 13 essentially proves this result. The main difference is that in Lemma 13 the proof was adapted to the specific schedule of the depths-$p$ epochs $(T_{k-1}^{(p)},T_{k}^{(p)}]$ for $k\in[N_{p}]$ for some $p\in[P]$. Within Proposition 9, because the epochs are also constructed online, we can replicate the same proof arguments with the online epochs $(T_{k-1},T_{k}]$ for $k\in[K]$.
Fix $w\geq 2$. We construct the equivalent alternative smooth process $(z_{a})_{a}$ together with probabilities $(\gamma_{a})_{a}$ as follows. On each epoch $k\in[K]$, we enumerate
$\left\{t\in(T_{k-1},T_{k}]:\gamma_{T_{k-1}}(t)\geq q\right\}=:\{t_{1}^{(k)}<\ldots<t_{b_{k}}^{(k)}\}.$
Using the same notations as in the proof of Lemma 13, we denote $\gamma_{l}^{(k)}:=\gamma_{T_{k-1}}(t_{l}^{(k)})$ for all $l\in[b_{k}]$. From now the construction of the alterative smooth process is identical. The length of the sequence is now $A=a_{K}$.
We now construct the functions $g_{a}$ for $a\in[T]$. As in the original proof we let $g_{a}=0$ for $a>A$. For $a\leq A$, letting $t_{l}^{(k)}$ be the time used to construct $z_{a}=x_{t_{l}^{(k)}}$, we let $f_{l}^{(k)},h_{l}^{(k)}$ such that
$\mathbb{P}\left(f_{l}^{(k)}(x_{t_{l}^{(k)}})\neq h_{l}^{(k)}(x_{t_{l}^{(k)}})\mid\mathcal{H}_{t_{l}^{(k)}-1}\right)\geq(1-\zeta)\gamma_{a},$
for some fixed value $\zeta>0$ then pose $g_{a}=\mathbbm{1}[f_{l}^{(k)}\neq h_{l}^{(k)}]$. Another difference with the proof of Lemma 13 is that we essentially have $\epsilon=0$-covers, which significantly simplifies the analysis. All the rest of the proof holds by using ~:=$\tilde{\mathcal{F}}:=\{\mathbbm{1}[f\neq g]:f,g\in\mathcal{F}\}$ instead of $\mathcal{F}_{p}$. Altogether, we obtain that with probability at least $1-\delta$,
$\sum_{a=1}^{A}\gamma_{a}$ $\lesssim\frac{\ln^{2}T}{q\sigma}\left(\ln\mathbb{E}_{\mu}\left[W_{8T\ln(\frac{T}{\delta})/\sigma}\left(\tilde{\mathcal{F}}\right)\right]+\ln\frac{T}{\delta}+w\right)$
$\lesssim\frac{\ln^{2}T}{q\sigma}\left(d\ln\left(\frac{T}{\sigma}\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}+w\right).$
where in the last inequality we use the fact that $\tilde{\mathcal{F}}$ has VC dimension at most $2d$ and Proposition 18 to bound the Wills functional. Furthering the bounds with the same arguments as in the proof of Lemma 13 and letting $w=w(T,\delta)=d\ln\left(\frac{T}{\sigma}\ln\frac{1}{\delta}\right)+\ln\frac{T}{\delta}+2\geq 2$ ends the proof.
For the bound in expectation, it suffices to take $w=w(T)\geq 2$ and use the high probability bound with $\delta=1/T$, which implies
$\mathbb{E}\left|\left\{k\in[K]:\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)\cdot\mathbbm{1}[\gamma_{T_{k-1}}(t)\geq q]\geq w(T,\delta)\right\}\right|\leq\delta T+C\frac{\ln^{2}T}{q\sigma}\lesssim\frac{\ln^{2}T}{q\sigma}.$
$\blacksquare$
We are now ready to prove the main regret bound for Cover.
Proof of Theorem 8  Again, this is a simplified version of the proof of Theorem 4. Fix $f^{\star}\in\mathcal{F}$. Instead of using Lemma 7, we can simply use the equivalent classical regret bound for the Hedge algorithm. Taking the union bound over all runs of Hedge on each epoch $k\in[K]$ and assuming that $K\leq T$, the regret decomposition Eq. 14 simply becomes with probability at least $1-\delta T$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))$ $\lesssim\sum_{k=1}^{K}\sqrt{(T_{k}-T_{k-1})d\ln T}+K\ln\frac{T}{\delta}+\sum_{k=1}^{K}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}\left(f_{k,S}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))$
$\lesssim\sqrt{KdT\ln T}+K\ln\frac{T}{\delta}+\sum_{k=1}^{K}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}\left(f_{k,S}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t})).$
where we denoted by $f_{k,S}$ the function from the cover constructed at the beginning of epoch $(T_{k-1},T_{k}]$ that had the same values as $f^{\star}$ on prior epoch queries (see line 3 of Algorithm 3). In the last inequality we use Jensen’s inequality. The exact same arguments as for bounding $\Lambda_{k}^{(p)}$ in Lemma 12 imply that with probability at least $1-\delta$,
$\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}\left(f_{k,S}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))\leq 2\sum_{t=T_{k-1}+1}^{T_{k}}\gamma_{T_{k-1}}(t)+3\ln\frac{T}{\delta},\quad k\in[K].$
From there the rest of the proof is essentially the same as for Theorem 4. As in Eq. 21, Proposition 9 together with the bound above implies that with probability at least $1-\delta$,
|$\left|\left\{k\in[K]:\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}\left(f_{k,S}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))\geq 5q(T_{k-1}-T_{k})\right\}\right|\leq C\frac{\ln^{2}T}{q\sigma}$
whenever $q\geq C^{\prime}\frac{w(T,\delta)}{L-1}$ where $C,C^{\prime}>0$ are some universal constants, $w(T,\delta)$ is as defined in Proposition 9, and $L=\max_{k\in[K]}T_{k}-T_{k-1}$ is the minimum length of a period. Note that because $K\leq T$, we have $L=\left\lceil T/K\right\rceil$. From there, as when bounding the terms $\Lambda_{k}^{(p)}$, we define
$q_{0}:=C_{1}\cdot\max\left(\frac{w(T,\delta)}{L-1},\frac{\ln^{3}T}{\sigma K}\right),$
where $C_{1}$ is a constant that may depend on the constants $C,C^{\prime}$ from above. Then, we obtain that on an event of probability at least $1-c\delta\ln T$ for some constant $c>0$, we have
$\sum_{k=1}^{K}\sum_{t=T_{k-1}+1}^{T_{k}}\ell_{t}\left(f_{k,S}(x_{t})\right)-\ell_{t}(f^{\star}(x_{t}))\leq(1+c)q_{0}T.$
This is the equivalent of Eq. 25. Altogether, we obtain that with probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))\lesssim\sqrt{KdT\ln T}+Kw(T,\delta)+\frac{\ln^{3}T}{\sigma K}T.$
We then take the value $K=\lfloor\ln T\cdot(T/d)^{1/3}\sigma^{-2/3}\rfloor$. Note that of $K\geq T$, the regret bound from Theorem 8 is immediate. This is also the case if $d/\sigma\gtrsim T$. Hence, from now we suppose that $K\leq T$ and $d/\sigma\leq T$. Then, we obtain with probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\ell_{t}(f^{\star}(x_{t}))\lesssim\ln^{2}T\left(\frac{dT^{2}}{\sigma}\right)^{1/3}+Kw(T,\delta).$
We then turn this oblivious regret guarantee into an adaptive regret guarantee using the same arguments as for Theorem 4 in Section 5.4. Altogether, we obtain that with probability at least $1-\delta$,
$\sum_{t=1}^{T}\ell_{t}(\hat{y}_{t})-\inf_{f\in\mathcal{F}}\sum_{t=1}^{T}\ell_{t}(f(x_{t}))\lesssim\ln^{2}T\left(\frac{dT^{2}}{\sigma}\right)^{1/3}+Kd\ln\frac{T}{\delta}\lesssim\ln T\left(\frac{dT^{2}}{\sigma}\right)^{1/3}\ln\frac{T}{\delta}.$
In the last inequality we used $d/\sigma\leq T$. This ends the proof of the theorem. $\blacksquare$
Appendix E Concentration inequalities and technical lemmas
We first state Freedman’s inequality [25] which gives tail probability bounds for martingales. The following statement is for instance taken from [11, Theorem 1] or [2, Lemma 9].
Lemma 19 (Freedman’s inequality).
Let $(Z_{t})_{t\in T}$ be a real-valued martingale difference sequence adapted to filtration $(\mathcal{F}_{t})_{t}$. If $|Z_{t}|\leq R$ almost surely, then for any $\eta\in(0,1/R)$ it holds that with probability at least $1-\delta$,
$\sum_{t=1}^{T}Z_{t}\leq\eta\sum_{t=1}^{T}\mathbb{E}[Z_{t}^{2}\mid\mathcal{F}_{t-1}]+\frac{\ln 1/\delta}{\eta}.$
For our purposes, we need strengthened versions of tools that were used in prior works on smoothed online learning. We start by giving a strengthened version of [15, Lemma 3].
Lemma 20.
Let $(x_{t})\subset\mathcal{X}$ be a sequence of random variables and let $g_{t}:\mathcal{X}\to[0,1]$ be a sequence of random functions adapted to a filtration $(\mathcal{H}_{t})_{t\geq 0}$ such that $g_{t}$ is $\mathcal{H}_{t-1}$-measurable and $x_{t}\mid(\mathcal{H}_{t-1},g_{t})$ is $\sigma$-smooth with respect to some measure $\mu$. Let $x_{s}^{\prime}$ be a tangent sequence. Finally, let $\tau$ be a stopping time for the filtration $(\mathcal{H}_{t})_{t\geq 0}$. Then,
$\sum_{t=1}^{\tau}\mathbb{E}[g_{t}(x_{t})\mid\mathcal{H}_{t-1},g_{t}]\leq 3\sqrt{\frac{\tau(1+2\ln\tau)}{\sigma}\left(1+\ln\tau+\sum_{t=1}^{\tau}\frac{1}{t}\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t-1},g_{t}]\right)}.$
As an important remark, compared to [15, Lemma 3], the bound from Lemma 20 has an improved dependency in $\sigma$. The bound is proportional $1/\sqrt{\sigma}$ instead of $1/\sigma$, which is needed to achieve the tight regret bounds from Theorem 4. Indeed, the lower bound from Theorem 5 also grows as $1/\sqrt{\sigma}$.
To prove Lemma 20 we first need to generalize [15, Lemma 2] as follows.
Lemma 21.
Let $(a_{t})_{t\in\mathbb{N}}$ be a sequence of numbers in $[0,1]$ such that $a_{0}>0$. For $K\geq 1$ and $T\geq 1$, define
$B_{T}(a,K):=\left\{t\in[T]:a_{t}\geq\frac{K}{t}\sum_{s=0}^{T}a_{s}\right\}.$
Then, for any $\epsilon\in(0,1]$, it holds that $|B_{T}(a,K)|\leq\epsilon T+\ln\frac{T}{a_{0}}$ for any $K\geq\frac{1}{\epsilon}\ln\frac{T}{a_{0}}$.
Proof  The proof is a simple adaptation from that of [15, Lemma 2], we only detail the modifications. As in the original proof, we define a new sequence $(b_{t})_{t\in\{0,\ldots,T\}}$ such that $b_{0}=a_{0}$ and for $t\in[T]$,
$b_{t}=\begin{cases}0&t\notin B_{T}(a,K)\\ \frac{K}{t}\sum_{s=0}^{t}b_{s}&t\in B_{T}(a,K).\end{cases}$
Their arguments show that $b_{t}\in[0,1]$ for all $t\in[T]$ and $B_{T}(a,K)=B_{T}(b,K)$ hence it suffices to focus on the sequence $b$. We enumerate $B_{T}(b,K)=\{t_{1}<\ldots<t_{i}\}\subset[T]$. Their arguments show that
$1\geq b_{t_{i}}=\frac{K}{t_{i}}\cdot\prod_{j=1}^{i-1}\left(1+\frac{K}{t_{j}}\right)b_{0}.$
We recall that $b_{0}=a_{0}$. Following their arguments, we obtain
$|B_{T}(a,K)|=i\leq\frac{\ln\frac{T}{Ka_{1}}}{\ln\left(1+\frac{K}{T}\right)}\leq\left(\frac{T}{K}+1\right)\ln\frac{T}{Ka_{1}}\leq\left(\frac{T}{K}+1\right)\ln\frac{T}{a_{1}},$
where in the second inequality we used $\ln(1+x)\geq\frac{x}{1+x}$ for all $x\geq 0$. This ends the proof. $\blacksquare$
We are now ready to prove Lemma 20. The proof is essentially the same as [15, Lemma 3], we give it for completeness.
Proof of Lemma 20  Using the same notations as in [15], let $p_{t}$ denote the law of $x_{t}$ conditioned on $\sigma(\mathcal{H}_{t-1},g_{t})$. By assumption, $\tau$ is a stopping, hence $\{\tau\geq t\}$ is $\mathcal{H}_{t-1}$-measurable. Then, denoting by $Z\sim\mu$ a random variable independent from $(x_{t},g_{t})_{t\geq 0}$ we have
$\sum_{t=1}^{\tau}\mathbb{E}[g_{t}(x_{t})\mid\mathcal{H}_{t-1},g_{t}]=\sum_{t=1}^{\tau}\mathbb{E}_{Z}\left[\frac{dp_{t}}{d\mu}(Z)g_{t}(Z)\mid p_{t},g_{t}\right]=\mathbb{E}_{Z,g_{t}}\left[\sum_{t=1}^{\tau}\frac{dp_{t}}{d\mu}(Z)g_{t}(Z)\mid\tau,g_{t},p_{t},t\leq\tau\right].$
Next, for any $K=\frac{1}{\epsilon}(1+\ln\frac{\tau}{\sigma})\geq 1$ where $\epsilon\in(0,1]$ will be specified later, we define $B_{\tau}(K)$ as in Lemma 21 to the sequence $(\sigma\frac{dp_{t}}{d\mu}(Z))_{t\in[\tau]}$ augmented with the value $a_{0}=\sigma$ at $t=0$. That is, we let
$B_{\tau}(K):=\left\{t\leq\tau:\frac{dp_{t}}{d\mu}(Z)\geq\frac{K}{t}\left(1+\sum_{s<t}\frac{dp_{s}}{d\mu}(Z)\right)\right\}.$
Note that because $(x_{t})_{t\in[T]}$ is $\sigma$-smooth, the constructed sequence has values in $[0,1]$. Then, Lemma 21 shows that $|B_{\tau}(K)|\leq\epsilon\tau+\ln\frac{T}{\sigma}$. Furthering the previous bounds and taking $K=2\ln(\tau)/\epsilon$, we then obtain
$\sum_{t=1}^{\tau}\frac{dp_{t}}{d\mu}(Z)g_{t}(Z)$ $\overset{(i)}{\leq}\frac{|B_{\tau}(K)|}{\sigma}+\sum_{t=1}^{\tau}\frac{K}{t}\left(1+\sum_{s=1}^{t-1}\frac{dp_{s}}{d\mu}(Z)g_{t}(Z)\right)$
$\leq\frac{\epsilon\tau+\ln\frac{\tau}{\sigma}}{\sigma}+\frac{1+\ln\frac{\tau}{\sigma}}{\epsilon}\left(1+\ln\tau+\sum_{t=1}^{\tau}\frac{1}{t}\sum_{s=1}^{t-1}\frac{dp_{s}}{d\mu}(Z)g_{t}(Z)\right).$ (45)
In $(i)$ we used the fact that $g_{t}$ has values in $[0,1]$ and that the process $(x_{t})_{t}$ is $\sigma$-smooth. The additional $1$ comes from the fact that $\tau\notin B_{\tau}(K)$. We take the value
$\epsilon=\sqrt{\frac{\sigma(1+2\ln\tau)}{\tau}\left(1+\ln\tau+\sum_{t=1}^{\tau}\frac{1}{t}\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t-1},g_{t}]\right)}.$
Note that if $\epsilon>1$, the bound from Lemma 20 is immediate since $\sigma\in(0,1]$ and we could have bounded the sum by $\tau$ directly. Similarly, if $\sigma\leq 1/\tau$ the bound is also immediate. Therefore, from now we suppose that $\epsilon\leq 1$ and $\sigma\geq 1/\tau$. In particular, this implies that $\epsilon\tau\geq\ln\tau$ Then, taking the expectation over $Z$ in Eq. 45 gives
$\sum_{t=1}^{\tau}\mathbb{E}[g_{t}(x_{t})\mid\mathcal{H}_{t-1},g_{t}]$ $\leq\frac{\epsilon\tau+2\ln\tau}{\sigma}+\frac{1+2\ln\tau}{\epsilon}\left(1+\ln\tau+\sum_{t=1}^{\tau}\frac{1}{t}\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t-1},g_{t}]\right)$
$\leq\frac{2\epsilon\tau}{\sigma}+\frac{1+2\ln\tau}{\epsilon}\left(1+\ln\tau+\sum_{t=1}^{\tau}\frac{1}{t}\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t-1},g_{t}]\right)$
$\leq 3\sqrt{\frac{\tau(1+2\ln\tau)}{\sigma}\left(1+\ln\tau+\sum_{t=1}^{\tau}\frac{1}{t}\sum_{s=1}^{t-1}\mathbb{E}[g_{t}(x_{s}^{\prime})\mid\mathcal{H}_{t-1},g_{t}]\right)}.$
This gives the desired result. $\blacksquare$
Next, we provide a high-probability version of [15, Theorem 2]. As a remark, this is only needed to obtain our high-probability oblivious regret bounds. In order to get expected oblivious regret bounds it suffices to use [15, Theorem 2] directly. This is however necessary to obtain our adaptive regret bounds in the case of function classes $\mathcal{F}$ with finite VC dimension, since these use the high-probability oblivious regret bounds to achieve low regret compared to a covering of the function class $\mathcal{F}$.
Lemma 22.
Let $\mathcal{F}:\mathcal{X}\to[0,1]$ be a function class and let $(x_{t})_{t\in[T]}$ be a smooth stochastic process with respect to some base measure $\mu$ on $\mathcal{X}$. Denote by $(x_{t}^{\prime})_{t\in[T]}$ a tangent sequence to $(x_{t})_{t\in[T]}$. Then, there exists a constant $C_{0}\geq 1$ such that for any $c>0$ and $\delta\in(0,1/2]$, with probability at least $1-\delta$,
$\sup_{f\in\mathcal{F}}\sum_{t=1}^{\tau}f(x_{t}^{\prime})-(1+2c)f(x_{t})\leq C_{0}\frac{(1+c)^{2}}{c}\left(\ln\mathbb{E}_{\mu}\left[W_{2T\ln(\frac{T}{\delta})/\sigma}\left(\frac{c}{1+c}\mathcal{F}\right)\right]+\frac{1}{c}\ln\frac{1}{\delta}\right).$
Note that compared to [15, Theorem 2], Lemma 22 bounds the sum of the values $f(x_{t})-(1+2c)f(x_{t}^{\prime})$ instead of $f^{2}(x_{t})-(1+2c)f^{2}(x_{t}^{\prime})$. Up to considering the function class $\mathcal{F}^{2}=\{f^{2}:f\in\mathcal{F}\}$, this implies the same result up to constants in light of [41, Theorem 4.1] which implies that for any $1$-Lipschitz real-valued function $\psi$, we have $W_{m}(\psi\circ\mathcal{F})\leq W_{m}(\mathcal{F})$ where $\psi\circ\mathcal{F}=\{\psi\circ f:f\in\mathcal{F}\}$.
Proof  We follow similar arguments as in the proof of [15, Theorem 2]. At the high level, the result is obtained by following the proof therein and turning each expectation step to a high-probability one. Using the same notations therein, their proof implies that the left hand side $\sup_{f\in\mathcal{F}}\sum_{t=1}^{\tau}f(x_{t}^{\prime})-(1+2c)f(x_{t})$ has the same distribution as
$\sup_{f\in\mathcal{F}}\sum_{t=1}^{\tau}(1+c)\epsilon_{t}\left(f(\boldsymbol{x}_{t}(\epsilon))-f(\boldsymbol{x}_{t}^{\prime}(\epsilon))\right)-c\left(f(\boldsymbol{x}_{t}(\epsilon))+f(\boldsymbol{x}_{t}^{\prime}(\epsilon))\right)\\ \leq\underbrace{\sup_{f\in\mathcal{F}}\left\{\sum_{t=1}^{\tau}(1+c)\epsilon_{t}f(\boldsymbol{x}_{t}(\epsilon))-cf(\boldsymbol{x}_{t}(\epsilon))\right\}}_{A}+\underbrace{\sup_{f\in\mathcal{F}}\left\{\sum_{t=1}^{\tau}-(1+c)\epsilon_{t}f(\boldsymbol{x}_{t}(\epsilon))-cf(\boldsymbol{x}_{t}(\epsilon))\right\}}_{A^{\prime}}.$
They then note that $A$ and $A^{\prime}$ have the same distribution by the symmetry of the Rademacher variables $\epsilon_{t}$, hence we can focus on bounding $A$ then use the union bound. Now introduce i.i.d. standard Gaussians $\xi_{1},\ldots,\xi_{T}$ independent from all other random variables. We also fix a function $\hat{f}\in\mathcal{F}$ such that
$\sum_{t=1}^{T}(1+c)\epsilon_{t}\hat{f}(\boldsymbol{x}_{t}(\epsilon))-c\hat{f}(\boldsymbol{x}_{t}(\epsilon))\geq(1-\eta)A,$
for some fixed parameter $\eta\in(0,1)$. Conditionally on other variables, including $\hat{f}$, the variables $|\xi_{1}|,\ldots,|\xi_{T}|$ are still i.i.d. and we note that $\sqrt{\frac{\pi}{2}}(1+c)\epsilon_{t}|\xi_{t}|\hat{f}(\boldsymbol{x}_{t}(\epsilon))$ is sub-Gaussian with parameter $C(1+c)^{2}\hat{f}^{4}(\boldsymbol{x}_{t}(\epsilon))$ for some universal constant $C\geq 1$. Applying the classical concentration bound for independent sub-Gaussian random variables, we obtain
$\sum_{t=1}^{T}\epsilon_{t}\left(\sqrt{\frac{\pi}{2}}|\xi_{t}|-1\right)\hat{f}(\boldsymbol{x}_{t}(\epsilon))\leq\sqrt{2C_{1}(1+c)^{2}\sum_{t=1}^{T}\hat{f}(\boldsymbol{x}_{t}(\epsilon))\cdot\ln\frac{1}{\delta}}.$ (46)
Here, we also used the fact that $\hat{f}$ takes values in $[0,1]$. Denote by $\mathcal{F}_{\delta}$ this event. We next consider the event
$\mathcal{G}_{\delta}:=\left\{\sum_{t=1}^{T}\hat{f}(\boldsymbol{x}_{t}(\epsilon))\leq\frac{8C_{1}(1+c)^{2}}{c^{2}}\ln\frac{1}{\delta}\right\}.$
Note that on the event $\mathcal{G}_{\delta}$, we directly have
$A\leq\frac{1}{1-\eta}\sum_{t=1}^{T}\hat{f}(\boldsymbol{x}_{t}(\epsilon))\leq\frac{8C_{1}(1+c)^{2}}{c^{2}(1-\eta)}\ln\frac{1}{\delta}$
On the other hand, on $\mathcal{F}_{\delta}\cap\mathcal{G}_{\delta}^{c}$, we can further bound Eq. 46 by $\frac{c}{2}\sum_{t=1}^{T}\hat{f}(\boldsymbol{x}_{t}(\epsilon))$. Then, we obtain
$A$ $\leq\frac{1}{1-\eta}\left(\sum_{t=1}^{T}(1+c)\epsilon_{t}\hat{f}(\boldsymbol{x}_{t}(\epsilon))-c\hat{f}(\boldsymbol{x}_{t}(\epsilon))\right)$
$\leq\frac{1}{1-\eta}\left(\sum_{t=1}^{T}\sqrt{\frac{\pi}{2}}(1+c)\epsilon_{t}|\xi_{t}|\hat{f}(\boldsymbol{x}_{t}(\epsilon))-\frac{c}{2}\hat{f}(\boldsymbol{x}_{t}(\epsilon))\right)$
$\leq\frac{1}{1-\eta}\left(\sum_{t=1}^{T}\sqrt{\frac{\pi}{2}}(1+c)\epsilon_{t}|\xi_{t}|\hat{f}(\boldsymbol{x}_{t}(\epsilon))-\frac{c}{2}\hat{f}^{2}(\boldsymbol{x}_{t}(\epsilon))\right)$
$\leq\frac{\pi(1+c)^{2}}{2c(1-\eta)}\underbrace{\sup_{f\in\mathcal{F}}\sum_{t=1}^{T}c^{\prime}\epsilon_{t}|\xi_{t}|f(\boldsymbol{x}_{t}(\epsilon))-\frac{c^{\prime 2}}{2}f^{2}(\boldsymbol{x}_{t}(\epsilon))}_{B},$
where $c^{\prime}=\sqrt{\frac{2}{\pi}}\frac{c}{1+c}$. In the third inequality we used the fact that the functions have values in $[0,1]$. Note that $\epsilon_{t}|\xi_{t}|$ has the same distribution as $\xi_{t}$. Hence defining $\boldsymbol{x}_{t}(\xi):=\boldsymbol{x}_{t}(\text{sign}(\xi))$, $B$ has the same distribution as if we replaced $\epsilon_{t}|\xi_{t}|$ by $\xi_{t}$, and replaced $\boldsymbol{x}_{t}(\epsilon)$ with $\boldsymbol{x}_{t}(\xi)$. Let $\mathcal{E}_{\delta}$ be the same event as in the proof of [15, Theorem 2] on which $\boldsymbol{x}_{t}(\epsilon)\in\{Z_{t,j},j\in[k]\}$ for all $t\in[T]$, where $k=\left\lceil\frac{1}{\sigma}\ln\frac{T}{\delta}\right\rceil$. We have $\mathbb{P}(\mathcal{E})\geq 1-Te^{-\sigma k}\geq 1-\delta$. Then, the arguments in [15, Theorem 2] show that
$\mathbb{E}\left[\exp\left(\mathbbm{1}[\mathcal{E}_{\delta}]\cdot B\right)\right]$ $\leq\mathbb{E}_{Z_{t,j}\sim\mu}W_{kT}\left(c^{\prime}\cdot\mathcal{F}\right).$
In particular, Markov’s inequality shows that with probability at least $1-\delta$,
$\mathbbm{1}[\mathcal{E}_{\delta}]\cdot B\leq\ln\mathbb{E}_{Z_{t,j}\sim\mu}W_{kT}\left(c^{\prime}\cdot\mathcal{F}\right)+\ln\frac{1}{\delta}.$
Denote by $\mathcal{H}_{\delta}$ this event. Putting everything together shows that on $\mathcal{E}_{\delta}\cap\mathcal{F}_{\delta}\cap\mathcal{H}_{\delta}$,
$A\leq\frac{8C_{1}(1+c)^{2}}{c^{2}(1-\eta)}\ln\frac{1}{\delta}+\frac{\pi(1+c)^{2}}{2c(1-\eta)}\left(\ln\mathbb{E}_{Z_{t,j}\sim\mu}W_{kT}\left(c^{\prime}\cdot\mathcal{F}\right)+\ln\frac{1}{\delta}\right),$
which has probability at least $1-3\delta$. We then use the union bound to similarly bound $A^{\prime}$. This shows that for some universal constant $C_{2}\geq 1$, with probability at least $1-6\delta$,
$\sup_{f\in\mathcal{F}}\sum_{t=1}^{T}f(x_{t})-(1+2c)f(x_{t}^{\prime})\leq C_{2}\left(\frac{(1+c)^{2}}{c}\ln\mathbb{E}_{Z_{t,j}\sim\mu}W_{2T\ln(\frac{T}{\delta})/\sigma}\left(c^{\prime}\cdot\mathcal{F}\right)+\frac{(1+c)^{2}}{c^{2}}\ln\frac{1}{\delta}\right).$
Noting that $c^{\prime}\leq\frac{c}{1+c}$, this gives the desired result. $\blacksquare$